### v1

# A Modified 3D Biconical Outflow Model: Spatial Constraints on AGN-driven Outflows

Abstract
We present a modified outflow model and its application to constrain ionized outflow properties of active galactic nuclei (AGNs). By adding a rotating disk component to the biconical outflow model of Bae16, we find that models with a rotating disk require faster launching velocities ($\lesssim$ 1500 km s-1) than outflow-only models to be consistent with the observed gas kinematics of local type 2 AGNs. We perform Monte Carlo simulations to reproduce the observed distribution of gas kinematics of a large sample ($\sim$ 39,000), constraining the launching velocity and opening angle. While the launching velocity is moderate for the majority of the local AGNs, the notable cases of 2 - 5 % show strong outflows with $V_{max}\sim 1000-1500$ km s-1. By examining the seeing effect based on the mock integral field unit data, we find that the outflow sizes measured based on velocity widths tend to be overestimated when the angular size of the outflow is comparable to or smaller than the seeing. This result highlights the need for more careful treatments of the seeing effect in the outflow size measurement, yet it still supports the lack of global feedback by gas outflows for local AGNs.
Active galactic nuclei (16), Galaxy kinematics (602), Galaxy winds (626)
software: astropy (Astropy22), matplotlib (matplotlib), numpy (numpy), scipy (scipy)
show]kcs1996kcs@snu.ac.kr
1 Introduction
Accreting supermassive black holes (SMBHs), or active galactic nuclei (AGNs), emit a vast amount of energy through various channels, i.e., radiation, gas outflows, and relativistic jets, which are thought to interact with the interstellar medium (ISM) and eventually affect star formation in host galaxies (e.g., Silk98; Fabian12). Recent theoretical studies commonly implement AGN feedback to reproduce primary observables, i.e., galaxy luminosity function (e.g., Hopkins08; Weinberger17; Wellons23). Nonetheless, the overall role and detailed mechanism of AGN feedback remains unclear in observations (see reviews by Harrison17; Harrison24).
AGN-driven gas outflows are widely considered as a primary candidate for a feedback channel. It is now well-known that these outflows are multiphase (e.g., Fiore17; Fluetsch19; Riffel23), ubiquitous (e.g., Mullaney13; Woo16; Rakshit18), and kpc-scales (e.g., Harrison14; Kim23). In particular, the broad wing component of the [O III] $\lambda 5007$ emission line is frequently utilized to investigate kpc-scale warm ionized gas outflows (Boroson05; Greene05). Large spectral surveys such as Sloan Digital Sky Survey (SDSS) enabled the scrutiny of integrated outflow properties within the central region for several tens of thousands of AGNs and their correlations with host galaxy properties (e.g., Woo17; Woo20).
On the other hand, using spatially resolved spectroscopy, many other studies have focused on the detailed outflow structures for a limited number of targets ($\lesssim$ several tens). One of the primary findings in early studies is that the ionized outflows are well-described by a biconical model with two axisymmetric cones (Crenshaw00a; Crenshaw00b; Veilleux01; Das05; Das06). Subsequent studies took into account the dust extinction by the galactic plane (e.g, Crenshaw10a; Crenshaw10b; Fischer10; Fischer11; Fischer13; Bae16) or a rotating disk (e.g., Muller-Sanchez11) to reproduce the [O III] line profile and diverse narrow line region (NLR) kinematics. Marconcini23 recently developed a new modeling tool for analyzing 3D gas kinematics based on the 3D distribution of numerous gas clouds. By using a sample of nearby AGNs, these studies have provided various constraints on the geometry, extent, and energetics of outflows, which are crucial for quantifying outflow energetics and their impact on the host galaxy evolution.
Despite the importance of estimating the outflow parameters, it is challenging to expand these works to a large sample due to observational limitations. For instance, spatially resolved spectroscopic observations are still expensive to be applied to a large sample. Furthermore, while integral field unit (IFU) surveys such as Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) or Sydney- AAO (Australian Astronomical Observatory) Multi-object Integral-field spectroscopy (SAMI) provide spatially resolved kinematics for $\lesssim$ thousands of AGNs (Wylezalek20; Deconto-Machado22; Gatto24; Oh24), they typically have a seeing-limited resolution (FWHM $\sim$ 1″- 2″), which is often insufficient to resolve outflow structures at moderate redshifts. For example, a few kpc scales of outflows at z $\sim$ 0.1 (1 kpc $\sim$ 0.5″) are barely resolved even under the best seeing conditions on the ground. While space telescopes such as Hubble Space Telescope (HST) and James Webb Space Telescope (JWST) or adaptive optics (AO) can improve the spatial resolution down to a diffraction-limit (FWHM $\sim$ 0.1″), the detailed geometry and kinematics can only be resolved in rare cases such as nearby AGNs (e.g., Fischer13; Polack24) or galactic scale outflows (e.g., Vayner21a; Wylezalek22; Cresci23).
Nevertheless, detailed outflow models and simulated mock data can be efficiently utilized to overcome the limited spatial resolution and constrain the physical parameters of the outflows, which are not directly measurable. Bae16 constructed a 3D biconical outflow model combined with a simple extinction by a thin dust plane and investigated the effect of each parameter on the gas kinematics. Their main findings are as follows: (1) velocity dispersion is higher for larger intrinsic outflow velocity or bicone inclination; (2) velocity offset increases with higher dust extinction; (3) a maximum velocity of 500 - 1000 km s-1can explain most of the observed gas kinematics of local SDSS type 2 AGNs from Woo16; and (4) higher intrinsic velocities tend to have wider bicone opening angles. However, this model did not account for rotating disk emission and seeing convolution, which are critical for reproducing 2D gas kinematics.
[…]
2 Modified 3D Biconical Outflow Model
We summarize the original outflow model from Bae16 in Section 2.1 and demonstrate three major modifications in Section 2.2, 2.3, and 2.4, respectively.
2.1 3D Biconical Outflow Model
Bae16 presented a 3D outflow model adopted from Crenshaw10a; Crenshaw10b, which consists of two axisymmetric outflow cones and a simple extinction factor by a thin galactic dust plane (see Figure 1). Once a set of 11 parameters is given, the model constructs 3D distributions of flux and velocity. By collapsing the 3D model along the line-of-sight (LOS), 2D distributions of flux, velocity, and velocity dispersion are generated. Bae16 also calculated flux-weighted velocity and velocity dispersion using the 2D maps for a direct comparison with the observed [O III] velocity and velocity dispersion (VVD) distribution of SDSS type 2 AGNs from Woo16.
The input parameters define the bicone structure, 3D distributions of flux and kinematics, and the orientations of the bicone and the dust plane. For the bicone structure, three parameters determine the outflow size ($R_{out}$) and inner and outer half opening angles ($\theta_{in}$ and $\theta_{out}$). As illustrated by the dark gray region in Figure 1, we model gas outflows using a hollow bicone. For simplicity, we set all models in this study to have an outer half opening angle that is 20 larger than the inner one ($\theta_{out}=\theta_{in}+20^{\circ}$).
The flux distribution is determined by e-folding radius ($\tau$) and dust extinction factor ($A$). Since we only focus on relative flux distributions, we adopt an arbitrary central flux value ($f_{0}$). We assume that the intrinsic 3D flux distribution follows an exponential form based on the distance from the center ($R$). Thus, the intrinsic 3D flux distribution can be written as $f(R)=f_{0}e^{-R/\tau}$. Following Bae16, we assume the e-folding radius is 20% of the outflow size ($\tau=R_{out}/5$), which gives approximately two orders of magnitude fainter flux at the outflow edge than the center. The dust extinction factor is multiplied when collapsing the 3D flux distributions to 2D flux maps. If a voxel is located behind the galactic dust plane along the LOS, its flux will be dimmed by the extinction factor in the 2D map. The opening angles and the relative inclination compared to the galactic plane decide which voxels are subject to dust extinction.
The 3D velocity distributions are determined by two parameters, i.e., maximum velocity ($V_{max}$) and velocity profile. There are three options for the velocity profile as a function of the distance from the center ($R$): (1) a linear increase case ($v(R)=V_{max}(R/R_{out})$); (2) a linear decrease case ($v(R)=V_{max}(1-R/R_{out})$); and (3) a constant velocity ($v(R)=V_{max}$). In all three cases, the velocity is directed radially outward from the center. Following Bae16, we adopt the linear decrease case for all models in this study. Note that the maximum velocity can be regarded as a launching velocity, as it represents the intrinsic outflow velocity at the center. This 3D velocity distribution is then collapsed to the 2D velocity and velocity dispersion maps in a flux-weighted manner as follows:
$V(x,y)=\frac{\int v_{p}(x,y,z)f(x,y,z)\,dz}{\int f(x,y,z)\,dz},$ (1)
$\sigma^{2}(x,y)=\frac{\int v^{2}_{p}(x,y,z)f(x,y,z)\,dz}{\int f(x,y,z)\,dz}-V^{2}(x,y)$ (2)
where $V(x,y)$ and $\sigma(x,y)$ are 2D distributions of velocity and velocity dispersion in the $(x,y)$ sky plane (see example 2D maps in the top row of Figure 2). $f(x,y,z)$ and $v_{p}(x,y,z)$ are 3D flux and projected velocity values at $(x,y,z)$ voxel. The LOS is defined as the z-axis in this coordinate system.
Lastly, the orientations of the model are determined by position angles (PAcone and PAdisk) and inclinations ($i_{cone}$ and $i_{disk}$) of the bicone and the dust plane (or disk in Section 2.2). The orientations of the bicone and the dust plane are independent of each other, allowing them to rotate freely on the sky plane. However, for simplicity, we fix the PAs of the bicone and dust plane to 0 and 90, respectively, meaning that the bicone axis and the major axis of the plane extend along the north-south and east-west directions, respectively. This is because our mock data analysis in Section 3 focuses on 3″-extracted spectra and radial profiles, which are little affected by PAs on the sky plane.
Using this biconical model, Bae16 tested the impact of each parameter on the velocity offset and dispersion, and the line profile of the [O III] emission line. Nevertheless, the model lacked several components crucial for comparison with observational data. For instance, it did not consider rotating gas in a galactic disk, which is thought to mainly contribute to the narrow/core component of the typical [O III] line profile. Furthermore, to accurately compare the spatial properties of the simulated outflows with observations, the seeing convolution should be taken into account. Therefore, we add a rotating gas component (Section 2.2) and the seeing effect (Section 2.3) to the Bae16 model. In addition, to improve mock data compatibility, we switch from a coordinate system normalized by the outflow size to a physical scale in arcsec units based on the redshift and outflow size parameters (Section 2.4).
Figure 1: A schematic diagram of a biconical model example of Bae16. SMBH is located at the center of this plot, and the dust plane is shown as a red line between two cones. The dark gray area represents where the outflowing gas exists (i.e., between inner and outer opening angles), while the light gray area shows a hollow region. The LOS direction is shown as a black arrow on the right.
2.2 Rotating Component
Although the outflow is a primary kinematic component in many AGNs, the gas kinematics in the NLR are more complicated due to the presence of other effects. Among them, gas following the gravitational potential of the host galaxy is often observed as a narrow emission line that has similar kinematics to stellar absorption lines ($\sigma_{[\mathrm{OIII}],narrow}\sim\sigma_{*}$ Nelson96; Greene05). For example, if there is no non-gravitational effect, i.e., outflows, the gas kinematics are dominated by a rotating disk component (e.g., Luo19). Given that a majority of AGNs in the VVD diagram (see Figure 4(a)) are located at relatively small [O III] velocity offset and dispersion, the disk component should be included to properly reproduce the observed distribution of the gas kinematics. Even for AGNs with strong outflows, a rotational pattern is often observed in IFU data (Kang18; Kim23). Thus, rotation due to the host galaxy gravity is required to properly reproduce the demography of the gas kinematics as well as the characteristics in the spatially resolved kinematics.
To generate disk models, we make use of the GalPak3D package (Galaxy Parameters and Kinematics, Bouche15), which was originally developed to model the morphokinematics of 3D observational data. Since our goal is to investigate how the rotating gas component influences the total gas kinematics, we employ only the basic models generated by the package without adjusting details. Consequently, all disk models in this study are based on the default models of GalPak3D. The radial and perpendicular flux profiles are set to exponential ($n=1$) and Gaussian, respectively, and the rotation curve follows a hyperbolic tan (tanh) profile.
We regulate two flux parameters and three kinematic parameters of the disk model. The inclination and PA of the disk are assumed to be the same as those of the dust plane in the outflow model. For the flux, the total disk flux within the SDSS fiber size (3″) is first determined by the disk-to-cone flux ratio parameter ($f_{disk}/f_{cone}$) and the integrated flux of the outflow. We then determine the disk flux distribution by assuming an exponential disk with a given half-light radius parameter ($r_{1/2}$).
For the gas kinematics, the maximum rotational velocity ($V_{max,rot}$) and the turnover radius ($r_{t}$) determine the velocity profile with a tanh function as follows:
$v(r)=V_{max,rot}\;\mathrm{tanh}(r/r_{t}),$ (3)
where $r$ is the radius on the galactic plane. In addition, we adjust the intrinsic velocity dispersion ($\sigma_{0}$), which corresponds to the dynamical hotness of the disk.
Once the five disk parameters and the disk orientation (PAdisk and $i_{disk}$) are given, GalPak3D computes flux, velocity, and velocity dispersion maps. We then sum the 2D flux maps of the outflow and disk models. For velocity and velocity dispersion maps, we combine them by flux-weighted means. In Figure 2, the second row shows mock 2D maps of the disk model for an example with $i_{disk}\sim 120^{\circ}$ at $z=0.05$. We assume a stellar mass of $\sim 10^{10.5}\;\rm{M}_{\odot}$, which corresponds to the following parameters: $V_{max,rot}=198$ km s-1, $r_{1/2}=5.2$ kpc, $\sigma_{0}\sim$ 117 km s-1, $r_{t}\sim$ 2.5 kpc. The third row of Figure 2 shows the combined maps of the outflow and disk models with $f_{disk}/f_{cone}=1$. These combined maps demonstrate that the biconical structure is properly embedded in the background disk model.
2.3 Seeing Effect
The beam smearing effect by atmospheric seeing has been pointed out as a significant obstacle in spatially resolved studies of AGNs (e.g., Husemann16; Villar-Martin16; Singha22). Given that the seeing effect is more pronounced at higher redshifts, we need to include it in our model to account for outflows in a majority of AGN host galaxies.
To reproduce more realistic 2D maps, we convolve the combined 2D maps with a Gaussian kernel as follows:
$G(x,y)=\frac{1}{2\pi\sigma^{2}_{g}}\mathrm{exp}(-\frac{x^{2}+y^{2}}{2\sigma^{2}_{g}})$ (4)
where $\sigma_{g}$ is a seeing size parameter ($\sigma_{g}=\mathrm{FWHM}/(2\sqrt{2\mathrm{ln}2})$). In this study, we adopt two cases of the seeing size, 0.8″  and 1.5″. The former is a typical seeing size of high resolution optical IFU data, while the latter is a median seeing size in SDSS. The convolved 2D maps are calculated as follows:
$F_{c}(x,y)=F(x,y)*G(x,y),$ (5)
$V_{c}(x,y)=\frac{(V(x,y)\times F(x,y))*G(x,y)}{F_{c}(x,y)},$ (6)
$\begin{split}&\sigma_{c}^{2}(x,y)\\ =&\frac{((V^{2}(x,y)+\sigma^{2}(x,y))\times F(x,y))*G(x,y)}{F_{c}(x,y)}\\ &-V^{2}_{c}(x,y)\end{split}$ (7)
where $F_{c}(x,y),V_{c}(x,y)$, and $\sigma_{c}(x,y)$ are convolved flux, velocity, and velocity dispersion, respectively. In the bottom row of Figure 2, we show an example of convolved 2D maps. The red horizontal bar on the upper right represents the applied seeing size in this example (FWHM =0.8=0.8″). The seeing convolution blurs the central bicone structure, so the final images are more like a circular shape in all 2D maps. We will discuss the seeing effect in Section 4.3 and 5.3, focusing on implications for the outflow size measurements in IFU data.
2.4 Mock Data Compatibility
We update our model to calculate 2D maps in an (x, y) coordinate system defined in arcsec units based on the input redshift and physical size of the outflow, while Bae16 model adopted relative scales normalized by the outflow size. This approach offers several advantages: (1) the outflow model uses the same pixel units as the disk model, making it convenient to combine the two models; (2) mock data can be directly compared with real observational data. Accordingly, we specify the mock data by pixel size and field-of-view (FoV) parameters, and all calculations are performed using this pixel scale from given redshift and physical length parameters ($R_{out}$, $r_{1/2}$, and $r_{t}$).
Figure 2: 2D map examples of the modified model. From left to right, three 2D maps in each row present flux, velocity, and velocity dispersion of the corresponding model. First row: Bae16 biconical model. This model has the same geometry as Figure 1. Second row: Disk model from GalPak3D. Third row: The bicone + disk combined model. Last row: The combined model after seeing convolution. Red dashed circles at the center and red bars on the top right show the FWHM seeing size. We also plot 3″-aperture size as gray circles, which is the fiber size of SDSS spectra.
3 Mock Data Analysis
In this Section, we describe our main analysis procedures based on mock data generated from the modified outflow model. First, we generate mock SDSS data extracted from the central 3″-aperture, and compare them with the observed [O III] kinematics of a large sample of SDSS type 2 AGNs in Section 3.1. In Section 3.2, we produce mock IFU data to analyze outflow sizes and 2D distributions of flux, velocity, and velocity dispersion.
3.1 Mock SDSS Data
Bae16 integrated the [O III] velocity offset and dispersion within the outflow radius and compared them with the observed VVD distribution of $\sim$39,000 SDSS type 2 AGNs from Woo16. Here, we use our modified model to reexamine the comparison with the observed VVD distribution.
To simulate SDSS spectra, we calculate the integrated velocity offset and dispersion ($v_{int}$, $\sigma_{int}$) within the central 3″-aperture shown as gray circles in Figure 2. In this example ($z=0.05$ and $R_{out}$=3 kpc), the projected size of the bicone exceeds the SDSS fiber region, so its outer part is not included in the mock SDSS data. While this rarely occurs for models with low redshifts and large outflow sizes, it provides a representative example of how modifications affect the mock data.
As we integrate the kinematics within an aperture twice as large as the FWHM of the seeing (1.5″), the seeing convolution does not significantly impact the VVD distributions. On the other hand, the rotating gas component influences the integrated velocity offset and dispersion since it has a velocity offset near zero and a narrow width with comparable flux to the outflow component. Thus, we mainly investigate the effect of the rotating gas component on the VVD distribution.
3.1.1 Model Grids for the Observed VVD
To investigate the effect of outflow properties on the observed kinematics, we overlay various model grids on the observed VVD diagram. We mainly investigate five parameters: the dust extinction factor ($A$), opening angles ($\theta_{out}$ and $\theta_{in}$), bicone inclination ($i_{cone}$), launching velocity ($V_{max}$), and disk-to-cone flux ratio ($f_{disk}/f_{cone}$). We fix the disk inclination as 60, the outflow size as 1 kpc, and the stellar mass as $10^{10.5}\;\mathrm{M}_{\odot}$.
We first compare the models with and without the rotating disk component. Then we adjust the bicone inclination while keeping the opening angle fixed, and vice versa. We change the bicone inclination from -40 to 0 and the outer half opening angle from 30 to 60 with a 10 step. For each model grid, we also adjust the dust extinction rate from 0 to 100 % (i.e., $A$ = 1 to 0) and launching velocities by 500, 1000, and 2000 km s-1.
3.1.2 Mock VVD from Monte Carlo Simulation
We perform Monte Carlo (MC) simulations and generate mock VVD distributions by assuming a probability distribution function (PDF) of each parameter. The primary purpose of these MC simulations is to identify the best PDFs of the launching velocity and opening angle parameters that reproduce the observed VVD distribution, and ultimately, to constrain these outflow properties for the type 2 AGN sample.
We first take PDFs of various parameters, such as redshift, stellar mass, disk-to-cone flux ratio, and extinction, from the observed SDSS type 2 AGNs (see Figure 3). The PDF of redshift (0.04 to 0.28 with 0.04 steps) is directly applied to the MC simulations. The observed disk-to-cone flux ratio is determined by the flux ratio between broad/wing and narrow/core components when the [O III] line is fitted with a double Gaussian. If the [O III] line is fitted with a single Gaussian, we classify the emission as either disk-only or outflow-only based on the difference between stellar and [O III] velocity dispersion. Specifically, when $|\sigma_{\mathrm{[OIII]}}-\sigma_{*}|$ is less than 30 km s-1, we classify this single Gaussian as disk-only emission, i.e., $f_{disk}/f_{cone}=\infty$, otherwise as outflow-only emission, i.e., $f_{disk}/f_{cone}=0$. We also estimate the PDF of the dust extinction factor by using observed $\alpha/H$$\beta$ with $R^{A}_{V}=4.5$ (Vogt13). Note that we restrict the dust extinction rate to smaller than 80% (or $A>0.2$) since a higher extinction rate results in too broad mock VVD distributions along the V[OIII] axis, which are inconsistent with the observed distribution. Furthermore, an extinction rate higher than 80% is an extreme case, corresponding to $A_{\lambda 5007}\gtrsim$ 1.75 mag.
The stellar mass is taken from the MPA-JHU catalog  [^1] https://wwwmpa.mpa-garching.mpg.de/SDSS/ and divided into five bins ($\rm{log}\;M_{*}=$9.7 to 11.3 with 0.4 step). We calculate the four remaining disk parameters, i.e., $r_{1/2}$ (Dutton11), $V_{max,rot}$ (Torres-Flores11), $\sigma_{0}$ (Oh22), and $r_{t}$ (Amorisco10), using empirical relations with stellar mass. Thus, once the stellar mass is selected among the five values, these four disk parameters are determined accordingly. Note that we restrict our sample to type 2 AGNs in the MC simulations due to the large uncertainties in stellar mass for type 1 AGNs.
For the outflow model, we first adopt random inclinations and fixed PAs (PAcone=0 and PAdisk=90) for both the bicone and dust plane. The inclination of the dust plane is given from 7.5 to 172.5 with 15 step, and its PDF is proportional to the solid angle of the normal vector integrated from 0 to 360 of PA. Since we only deal with type 2 AGNs, the bicone inclination is restricted to less than 40, i.e., 7.5, 22.5, and 37.5, and its PDF is also proportional to the solid angle of the bicone axis. Then we assume that the PDF for five outflow size values (0.3, 1, 3, 6, 9 kpc) follows a Gaussian distribution with a mean of zero and a standard deviation of 2 kpc. Note that the outflow size has little effect on the kinematics in mock SDSS data.
Our main goal is to constrain two parameters, i.e., the launching velocity and opening angle of the outflow model. For these two parameters, we assume that their PDFs follow a truncated Gaussian function. We first set the range of the launching velocity to $100,\;\infty)$ km s-1  and restrict the outer opening angle values from 30 to 70 with a 10 step. We then run multiple MC simulations by changing the mean and standard deviation of the Gaussian to find the best $V_{max}$ and $\theta_{out}$ distributions that generate the best-matching mock VVD with the observed VVD. Each MC simulation generates 100,000 pairs of velocity offset and dispersion.
3.1.3 Kullback-Leibler Divergence and $D_{max}$
To quantitatively compare the mock VVDs with the observed VVD, we apply two statistical quantities that assess proximity between two 2D distributions. We first calculate Kullback-Leibler divergence ($D_{KL}(P||Q)$, KLDIV) as follows:
$D_{KL}(P||Q)=\sum P(v_{obs},\sigma_{obs})\;\mathrm{log}\frac{P(v_{obs},\sigma_{obs})}{Q(v_{int},\sigma_{int})},$ (8)
where $P(v_{obs},\sigma_{obs})$ and $Q(v_{int},\sigma_{int})$ are probability distributions of the observed and mock VVD, respectively. In brief, the Kullback-Leibler divergence is a measure of a statistical distance between two PDFs derived from information theory, i.e., the smaller the Kullback-Leibler divergence is, the more similar the two PDFs are.
The second quantity is the $D_{max}$ from the 2D Kolmogorov-Smirnov test, calculated from the public code ndtest [^2] Written by Zhaozhou Li, https://github.com/syrte/ndtest. Like the 1D Kolmogorov-Smirnov test, it calculates the maximum absolute difference between two cumulative distribution functions ($D_{max}$) and performs a p-value test of whether two PDFs originate from the same distribution. Although all p-values are too small to conclude that the observed and mock VVDs are identical distributions, we utilize the $D_{max}$ value to compare the relative superiority between the mock VVDs.
Note that we add random error smoothing to all mock VVDs based on typical SDSS uncertainties of velocity offset and dispersion to account for the observational uncertainties. To remove variations induced by this error smoothing, we repeat this process 100 times in each grid of $V_{max}$ and $\theta_{out}$ PDFs and take the mean and standard deviation of Kullback-Leibler divergence and $D_{max}$ values. If the mean Kullback-Leibler divergence (or $D_{max}$) for given parameter distributions is notably smaller than for the others, by more than its standard deviation, then that mock VVD distribution is considered a better reproduction of the observed data.
Figure 3: Histograms of physical properties of Woo16 type 2 AGNs. From left to right, each histogram shows redshift, stellar mass, flux ratio between disk and outflow components, and extinction factor based on Balmer decrement.
3.2 Mock IFU Data
We simulate mock 2D maps using various redshifts and outflow sizes to test the validity of previous methods for outflow size measurements. In particular, we focus on the dependence of the measured sizes on the seeing size normalized by the angular size of the outflow.
In our previous IFU studies, we developed a kinematic approach to measure the outflow size and applied it to AGNs at $z<0.3$ (Karouzos16a; Kang18; Bae17; Luo21; Kim23). The outflow region is defined as where the [O III] velocity dispersion is more enhanced than stellar velocity dispersion, which traces the gravitational potential of the host galaxies. Specifically, we first plot a radial profile of ratios between the [O III] and stellar velocity dispersion ($\sigma_{\mathrm{[OIII]}}/\sigma_{*}$) and define the outflow size as a radius where the ratio becomes unity. Then we quadratically subtract the seeing size ($\sigma_{g}$) from the outflow size to correct the seeing effect.
We apply the same procedure to mock IFU data and compare the measured outflow size with the input outflow size. Since we do not have stellar velocity dispersion in our model, we alternatively adopt the [O III] velocity dispersion extracted within a 3″-aperture of the convolved disk-only maps. This velocity dispersion of the disk model mimics the [O III] narrow/core component, which is an optimal proxy of stellar velocity dispersion (e.g., Greene05).
Another widely used method for outflow size measurements is W80-based method (e.g., Harrison14; Wylezalek20; Ruschel-Dutra21), where W80 is defined as a [O III] line width including 80% of the total line flux. Similar to the $\sigma_{\mathrm{[OIII]}}$-based method, the W80-based method defines the outflow region where the W80 is broader than a certain criterion, frequently set to 600 km s-1. Thus, the outflow size is determined as the radius of the furthest pixel from the center with W80 larger than 600 km s-1.
[…]
4 Result
4.1 Location of Model Grids in VVD Diagram
Figure 4: Model grids from the modified outflow model. Each color indicates a different launching velocity, 500 km s-1(cyan), 1000 km s-1(yellow), and 2000 km s-1(red): (a) The observed [O III] VVD distribution of type 2 AGNs (Woo16). This color map is also plotted as contours and gray color maps in the other three panels; (b) Model grids with (solid lines) and without (dashed lines) the disk. Each line shows the change in kinematics depending on extinction at a fixed inclination; (c) Model grids showing the effects of bicone inclinations and extinction rates when the disk component is included. Solid lines indicate variations with extinction at a fixed bicone inclination, while dotted lines indicate variations with inclination at a fixed extinction rate; (d) Model grids with different opening angles and extinction rates. Each solid and dotted line indicates the same bicone opening angle and extinction rate, respectively.
We investigate the dependency of outflow parameters on the location in the VVD diagram by comparing model grids with observations. The observed VVD distribution of local type 2 AGNs in Figure 4(a) is characterized by a V-shaped envelope and more AGNs with blueshifted [O III] than redshifted (Woo16).
First, we examine the effect of the rotating disk emission by comparing model grids with/without the disk component (Figure 4(b)), finding that both the integrated velocity offset and dispersion ($v_{int}$ and $\sigma_{int}$) decrease by including a disk component. This is because the disk component usually has a smaller velocity offset and dispersion than the outflow, resulting in smaller velocity offset and dispersion values in the combined model than in the outflow-only model. Thus, the modified model with a disk component requires a higher launching velocity ($\lesssim$ 1500 km s-1) to reproduce a similar VVD range, while most of AGNs in the VVD diagram were explained by $V_{max}\lesssim 1000$ km s-1(orange dashed lines) in the original model (Bae16).
Second, we revisit the effect of the bicone inclination and the dust extinction (Figure 4(c)). We find that both $\sigma_{int}$ and $v_{int}$ increase at a more inclined bicone, which is consistent with Bae16. For higher extinction, $\sigma_{int}$ decreases while $v_{int}$ increases because of the less contribution of the receding cone. Notably, the angle of the V-shape becomes smaller than that of the outflow-only model since the ratio of $|v_{int}|/\sigma_{int}$ decreases due to the zero velocity offset of the disk component.
Lastly, we investigate the effect of the opening angle (Figure 4(d)). We find that a wider opening angle generates a larger $\sigma_{int}$ due to a broader range of LOS velocities, while $v_{int}$ is almost constant regardless of the opening angle. Overall, the effect of the opening angle is consistent with Bae16.
4.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
We constrain the distributions of $V_{max}$ and $\theta_{out}$ parameters by comparing mock VVD distributions from the MC simulations with the observed VVD distribution. We conduct 315 realizations of MC simulations using Gaussian distributions, for which the mean and standard deviation of $V_{max}$ change from 150 to 450 km s-1and from 250 to 450 km s-1, respectively, in 50 km s-1steps. Mean and standard deviation of the $\theta_{out}$ parameter are set to 30, 40, 50 and 5, 10, 20, respectively. Note that the adopted mean and standard deviation are different from those of the resulting PDFs, as we are using a truncated Gaussian distribution.
Figure 5: Left: Kullback-Leibler divergence values depending on the PDF of the launching velocity. The PDF of the opening angle is fixed to the best. The mean and standard deviation of $V_{max}$ first determine the Gaussian distribution, and we use the truncated Gaussian from 100 km s-1as a PDF of the $V_{max}$. The minimum Kullback-Leibler divergence is found when the mean and standard deviation are 250 and 400 km s-1. Right: The same plot but $D_{max}$ from the 2D KS test. The minimum value is found when the mean and standard deviation are 350 and 300 km s-1.
Figure 6: Left: Kullback-Leibler divergence values depending on the PDF of the opening angle. The minimum Kullback-Leibler divergence is found when the mean and standard deviation are 30 and 10, respectively. Right: The same plot but $D_{max}$ from the 2D KS test. The minimum value is found when the mean and standard deviation are 30 and 5.
We first investigate variations of Kullback-Leibler divergence and $D_{max}$ values depending on the input launching velocity and outer opening angle PDFs in Figure 5 and 6. In each map, we fix the PDF of the other parameter to the best one. The Kullback-Leibler divergence and $D_{max}$ maps of the launching velocity show local minima near the center, ruling out the lower left and upper right corners. This indicates that the observed VVD distribution prefers a moderate range of launching velocity in general. On the other hand, the PDF of the opening angle prefers small values (i.e., lower left corner regions in Figure 6). Specifically, Kullback-Leibler divergence is minimized when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(250 km s-1, 400 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 10). $D_{max}$ exhibits a minimum value when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(350 km s-1, 300 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 5).
We present the best mock VVD distributions based on Kullback-Leibler divergence (left) and $D_{max}$ (right) in Figure 7. The two best mock VVD distributions are generally consistent with the observed one, reproducing both the V-shape feature and the overall range of the V[OIII] and $\sigma_{\mathrm{[OIII]}}$. The best mock VVD distribution based on Kullback-Leibler divergence exhibits a slightly larger $\sigma_{\mathrm{[OIII]}}$ range than $D_{max}$. Despite the extinction limit ($A\geq 0.2$), the V-shape envelope of the mock VVD distribution is slightly broader than the observed one, especially in the low $\sigma_{\mathrm{[OIII]}}$ regime. This may result from several factors, such as large observational uncertainties for weak outflow AGNs.
Figure 7: The best mock VVD distributions based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). The contours from the observed VVD distribution are overlaid. Dashed box indicates a selection criterion of strong outflow AGNs ($|\mathrm{V}_{\mathrm{[OIII]}}|>300$ km s-1or $\sigma_{\mathrm{[OIII]}}>400$ km s-1).
Figure 8: The best PDFs of the launching velocity based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). The histograms show the best PDFs for the entire mock VVD distributions, which are derived from Gaussian distributions truncated from 100 km s-1. The mean and standard deviation of the Gaussian are denoted in each panel.
Figure 9: The launching velocity distributions for a subset of strong outflow targets (outside the dashed boxes in Figure 7). The distributions are obtained from the best mock VVD distributions based on Kullback-Leibler divergence (left) and $D_{max}$ (right).
Figure 10: The best PDFs of the opening angle based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). Both follow Gaussian distributions truncated from 30 to 70.
The histograms in Figure 8 show the best PDFs of the launching velocity based on Kullback-Leibler divergence (left) and $D_{max}$ (right). The two quantities adopt similar best PDFs for the launching velocity. As predicted in Section 4.1, the launching velocities are mostly smaller than 1000 km s-1, while 2 - 5% exhibit higher launching velocities (1000 - 1500 km s-1).
In contrast, strong outflow AGNs exhibit significantly different $V_{max}$ distribution from the total sample. The histograms in Figure 9 represent $V_{max}$ distributions of strong outflow AGNs, which are located outside the dashed boxes in Figure 7 ($|\mathrm{V}_{\mathrm{[OIII]}}|>300$ km s-1or $\sigma_{\mathrm{[OIII]}}>400$ km s-1). Compared to the total distributions, strong outflow cases have fast launching velocities, mostly 800 - 1500 km s-1. Since these AGNs are primary targets in spatially resolved studies, this result constrains a typical range of their launching velocities.
Lastly, the best PDFs of the opening angle (Figure 10) are slightly different depending on the statistical quantities. Specifically, $\theta_{out}\geq 50^{\circ}$ is only exhibited in the best PDF based on Kullback-Leibler divergence, while the best PDF based on $D_{max}$ is concentrated on $\theta_{out}=30^{\circ}-40^{\circ}$. Nevertheless, the two best PDFs consistently favor relatively narrow outer opening angles, indicating that the opening angle of the local type 2 AGN is likely to be narrow, mostly 30 - 40.
4.3 Testing Outflow Sizes
Table 1: Model Parameters for Mock IFU Analysis
z $R_{out}$ $V_{max}$ $\theta_{out}$ $i_{cone}$ $i_{disk}$ AA $f_{disk}/f_{cone}$ $r_{t}$ $\sigma_{0}$ $V_{max,rot}$ $r_{1/2}$
[1] [2] [3] [4] [5] [6] [7] [8] [9] [10] [11] [12]
0.05 vs 0.3 1 kpc 1250 km s-1 40 25 120 0.2 1 2.5 kpc 117 km s-1 198 km s-1 5.2 kpc
Note. — (1) Redshift; (2) outflow size; (3) maximum (launching) velocity; (4) outer opening angle; (5) bicone inclination; (6) disk (and dust plane) inclination; (7) extinction factor (smaller value means more extinction); (8) flux ratio between disk and bicone components; (9) velocity turnover radius (disk); (10) intrinsic velocity dispersion (disk); (11) maximum rotational velocity (disk); (12) effective radius.
Figure 11: An example of mock IFU analysis when redshift is 0.05. Left: Radial profiles of $\sigma_{\mathrm{[OIII]}}$ (yellow markers) and W80 (blue markers). Large colored markers show the means of velocity widths in individual radial bins. Blue and yellow horizontal lines represent the criteria of the outflow region for W80 (600 km s-1) and $\sigma_{\mathrm{[OIII]}}$ ($\sigma_{*}$), respectively. Vertical dashed lines indicate half-width-half-maximum (HWHM) seeing size (red), input outflow size (black), the W80-based size (blue), and the $\sigma_{\mathrm{[OIII]}}$-based size (yellow). Right: 2D maps of velocity dispersion and W80. The colors of dashed line circles follow the same color scheme as vertical lines in the left panel.
Figure 12: The same plots as Figure 11 but when redshift is 0.3.
We investigate the validity of the outflow size measurements based on mock IFU data by testing the effect of seeing. We set two examples of the same physical properties except for their redshifts (0.05 vs 0.3, see Table 1) to contrast the angular sizes of the outflows by a factor of $\sim$ five for a given seeing (FWHM = 0.8″).
We present the spatial distributions of velocity widths of the two examples in Figure 11 and 12. In our previous IFU studies, radial profiles of velocity dispersion often exhibit constant values in the central region (i.e., initial plateau) before a smooth decrease (e.g., Karouzos16a; Kang18). This initial plateau is well reproduced in our mock IFU data, for both $\sigma_{\mathrm{[OIII]}}$ (yellow) and W80 (blue). Moreover, the radius of the initial plateau is comparable to the HWHM of the seeing (red dashed line), indicating that it mainly results from the seeing effect.
In the low redshift case (Figure 11), since the angular outflow size is larger than the seeing by more than a factor of two, the measured outflow size is consistent with the input value, indicating that the outflow size is well recovered by applying the velocity-width-based methods. In contrast, in the high redshift case (Figure 12), the angular outflow size becomes smaller than the seeing, leading to a significant overestimation of the outflow size even after quadratically subtracting the seeing size. Note that the measured size is also several times larger than the HWHM of the seeing size, as the seeing profile does not converge to zero even outside the FWHM. For example, if an outflow is unresolved, its seeing-convolved flux will follow the seeing profile. Thus, the outflow flux broadens the line profile at a radius further than the HWHM of the seeing, although its emission is dimmed by more than half compared to the center.
These results indicate that the seeing effect has to be carefully considered for interpreting IFU observations, i.e., initial plateau and outflow size. While the outflow size measured based on the emission line width is reliable when the seeing size is relatively small, the outflow size can be overestimated when the seeing size is similar to or larger than the angular size of the outflow.
5 Discussion
5.1 Comparison with Other Outflow Models
In this section, we discuss the characteristics of our model in comparison with the previous models developed over the last two decades. First, early outflow models did not include a rotating disk. For example, Crenshaw00a; Crenshaw00b constructed a hollow bicone model to reproduce position-velocity (PV) diagrams of nearby AGNs (e.g., Das05; Das06). Subsequently, Crenshaw10a; Crenshaw10b added a dust extinction component to the model to apply for dozens of nearby AGNs (e.g., Fischer10; Fischer11; Fischer13; Meena21; Falcone24). Since these models were developed for the HST observations of nearby AGNs, it is not essential to include the effect of the point spread function (PSF).
On the other hand, Carniani15 developed a model with a single approaching cone for testing their spectroastrometric analysis of powerful outflows in quasars at cosmic noon. Since the seeing effect is substantial for their high-z AGNs, they convolved each velocity channel map with the PSF, reproducing consistent [O III] flux, velocity, and velocity dispersion maps with the observations. Although their model lacks a disk component, the contribution of the disk component is presumably weak in these extremely luminous AGNs.
Second, other models were also reported for accommodating various kinematical considerations. For example, Muller-Sanchez11 constructed a model to analyze the coronal line regions (CLRs) using AO-assisted near-infrared IFU data. By combining a bicone model with a rotating disk, they successfully fitted CLR kinematics at the central several hundred pc, while no seeing effect was included.
Recently, Marconcini23 developed a new model (MOKA3D) based on Venturi17 by constructing a 3D distribution of point-like clouds to reproduce the observed IFU data. The model also convolves the 3D data and assigns a proper weight to each cloud. Since outflows often exhibit clumpy regions within their bicone geometry (e.g., Kakkad23), their model has a major advantage in reproducing various substructures, showing excellent agreements with the observations of both nearby and high-z AGNs (Cresci23; Marconcini23; Marconcini25a; Marconcini25b).
Compared to the previous models, which are mainly focused on fitting 3D data cubes of individual AGNs, our model is better suited for investigating the distributions of outflow parameters for a large sample, including the effect of seeing size and the disk component. Note that our model can also be utilized for reproducing the observation of individual AGNs (e.g., Shin19). Thus, our analysis with the modified model is complementary to the previous outflow models.
5.2 Constraints on the AGN-driven outflows
The launching velocity is the key parameter for quantifying outflow energetics (see Section 5.4), while the opening angle is related to how strongly gas outflows interact with the ISM. Thus, it is essential to constrain the launching velocity and opening angle of outflows for a large sample to investigate AGN feedback.
We report launching velocities of $\lesssim$ 1500 km s-1  and typical opening angles of 30-40 (Section 4.2). For a consistency check, we discuss the previous estimations of these two parameters, mostly from nearby AGNs. For example, Crenshaw00a; Crenshaw00b reported the inner/outer opening angles as 26/40 and the maximum outflow velocity as 1300 km s-1for NGC 1068, and inner/outer opening angles as 20/36 and the maximum outflow velocity as 750 km s-1for NGC 4151 (see also Das05; Das06). After including dust extinction, the outer opening angles are slightly changed to 40 for NGC 1068 and 33 for NGC 4151 (Crenshaw10b).
Their following studies also reported outflow parameters of other nearby AGNs (Crenshaw10a; Fischer10; Fischer11; Meena21; Falcone24). In particular, among the cases of 17 Seyfert galaxies investigated by Fischer13, the maximum velocity agrees with our range of $V_{max}\leq 1500$km s-1, except for NGC 1068. On the other hand, the outer opening angle distribution is slightly different. While a majority of their sample shows consistent opening angles (<<50) with our results, 7 AGNs have wider opening angles. The difference can be caused by a couple of effects. First, AGNs with higher luminosity and stronger outflows are likely to have wider opening angles (Bae16). For example, only 13% of our type 2 AGN sample have bolometric luminosity higher than $10^{44.5}$ erg/s, while more than half of their sample (9/17) are more luminous than $10^{44.5}$ erg/s. Second, their sample was selected based on clear signs of extended outflows, whereas we do not apply any criterion on the outflow extent. The inclusion of the disk component in our model can also induce differences.
[…]
The maximum velocity and the outer opening angle estimated by Muller-Sanchez11; Muller-Sanchez16 are also consistent with our constraints. On the other hand, for the best-fit parameters from Nevin18, while the maximum velocities ($V_{max}=230-720$ km s-1) also agree with our study, most of the opening angle values are much higher than the typical range of our constraint. This discrepancy also appears to be due to different biconical model and/or the characteristics of their sample (i.e., AGNs with double-peaked narrow emission lines). Other studies that constrained the opening angles of individual galaxies (e.g., Baron18; Riffel21; Costa-Souza24) are in agreement with our constraint.
5.3 True Outflow Size
[…]
In Section 4.3, we verified that the outflow size measured based on velocity widths can be significantly overestimated by the seeing, especially when a true outflow size is comparable to or smaller than the seeing size. To properly constrain the spatial scale of outflows, it is essential to investigate the seeing-limited cases in detail.
Previous studies reported notable spatial features in luminous type 2 AGNs, i.e., circularly extended NLR morphologies or enhanced W80 values with flat radial profiles (e.g., Liu13a; Liu13b; Harrison14). Interestingly, our model reproduces a similar flat W80 profile when the outflow is compact and the disk emission is much weaker than the outflow. Seeing convolution of a compact outflow generates a broad wing component with an almost constant velocity width across the entire FoV, while the disk emission is too faint to affect the velocity width, even at pixels far from the center. Thus, given that their samples consist of luminous type 2 quasars, implying relatively weak disk emission, the flat W80 profiles may indicate smaller outflow sizes.
Furthermore, Villar-Martin16 found that only 2 out of 15 AGNs in their sample show clearly resolved outflows by comparing the spatial profiles of seeing with the [O III] wing emission. Based on spectroastrometry and seeing sizes, they constrained lower and upper limits of the outflow sizes as several hundred pc and a few kpc, respectively. These results are in agreement with our constraints and represent observational evidence of the seeing effect in local AGNs.
[…]
In our previous IFU studies (Karouzos16a; Bae17; Kang18; Luo21; Kim23), the outflow sizes were determined based on the $\sigma_{\mathrm{[OIII]}}$-based method. Our main conclusion was that the outflow sizes are mostly less than several kpc scales, indicating a lack of galactic scale feedback. We also reported a kinematic outflow size–luminosity relation between the outflow size and the [O III] luminosity with a slope of $\sim$0.3. For this work, we subtracted the HWHM of the seeing size from the outflow size in quadrature to correct for the seeing effect. However, this correction may not be sufficient as shown in Section 4.3. Thus, our measurements can be conservatively considered as upper limits, particularly for higher redshift targets among our sample ($z>0.1$).
To test the reliability of the size measurements, we apply the $\sigma_{\mathrm{[OIII]}}$-based method to various mock data. For this test, we set the redshift z=0.1 as a representative case and restrict $V_{max}$ from 750 to 2000 km s-1 to focus on strong outflows (see Section 4.2). In Figure 13, we present the measured outflow size as a function of the input (true) size. As expected, the outflow size is overestimated by more than a factor of two when the true outflow size is smaller than the FWHM of the seeing ($\sim$1.52 kpc at z = 0.1). On the other hand, when the seeing is smaller than the true size, the measured size is generally consistent with the input size with slight underestimations (<< 0.3 dex). Therefore, the $\sigma_{\mathrm{[OIII]}}$-based method provides a good size estimation within 0.3 dex, or at least can be used as a reliable upper limit, depending on the relative seeing size.
We emphasize that although the $\sigma_{\mathrm{[OIII]}}$-based method overestimates the sizes of compact outflows, the main conclusion of Kim23, i.e., lack of global feedback, remains valid. Since the true size can be more compact than the measured one, the upper limit of the size still provides strong constraints on the spatial scales of AGN feedback.
It is also noteworthy that the measured sizes are broadly consistent within $\sim$ 0.3 dex at a given input size, regardless of other input parameters. There are some outliers with small measured sizes, which mostly have slow launching velocities (<< 1000 km s-1). Thus, if the launching velocity is fast enough, the $\sigma_{\mathrm{[OIII]}}$-based size has no significant dependence on other parameters.
5.4 Outflow Energetics
[…]
We estimate mass outflow rates ($\dot{\mathrm{M}}_{\mathrm{out}}$), momentum rates ($\dot{\mathrm{P}}_{\mathrm{out}}$), and kinetic powers ($\dot{\mathrm{E}}_{\mathrm{out}}$) of local type 2 AGNs to investigate feedback efficiency. These outflow rates can be determined for conical outflows as (e.g., Maiolino12; Fiore17)
$\dot{\mathrm{M}}_{\mathrm{out}}=3\;\mathrm{M}_{\mathrm{out}}\frac{v_{\mathrm{out}}}{\mathrm{R}_{\mathrm{out}}}$ (9)
$\dot{\mathrm{P}}_{\mathrm{out}}=\dot{\mathrm{M}}_{\mathrm{out}}v_{\mathrm{out}}$ (10)
$\dot{\mathrm{E}}_{\mathrm{out}}=\frac{1}{2}\dot{\mathrm{M}}_{\mathrm{out}}v^{2}_{\mathrm{out}}$ (11)
where $\mathrm{M}_{\mathrm{out}}$, $v_{\mathrm{out}}$, and $\mathrm{R}_{\mathrm{out}}$ are ionized gas mass, outflow velocity, and outflow size, respectively. We estimate ionized gas mass from the extinction-corrected [O III] luminosity, using Eq. 29 in Veilleux20 and a fixed electron density of 1000 cm-3.
In the case of outflow velocity, various definitions were adopted by previous studies, i.e., W80 (e.g., Wylezalek20; Ruschel-Dutra21), $\sqrt{v^{2}_{\mathrm{[OIII]}}+\sigma^{2}_{\mathrm{[OIII]}}}$ (e.g., Rakshit18; Kim23), or $v_{\mathrm{broad}}+2\sigma_{\mathrm{broad}}$ (e.g., Rupke13; Fiore17) to represent the flux-weighted mean outflow velocity. We derive the mass-weighted mean outflow velocity, using the radial profiles of outflow velocity $v(R)=V_{max}(1-R/R_{out})$ and flux $f(R)=f_{0}e^{-5R/R_{out}}$ in our model and assuming an electron density profile, $n_{e}(R)$ $\propto$ $1/R^{2}$. Since the mass is proportional to L/[$_{\mathrm{[OIII]}}/n_{e}$, the mass-weighted mean velocity is derived as:
$\begin{split}v_{\mathrm{out}}&=\frac{\int^{R_{out}}_{0}v(R)\mathrm{M}_{\mathrm{out}}(R)\,R^{2}dR}{\int^{R_{out}}_{0}\mathrm{M}_{\mathrm{out}}(R)\,R^{2}dR}\\ &=\frac{\int^{R_{out}}_{0}V_{max}(1-\frac{R}{R_{out}})\times f_{0}e^{-\frac{5R}{R_{out}}}n^{-1}_{e}(R)R^{2}\,dR}{\int^{R_{out}}_{0}f_{0}e^{-\frac{5R}{R_{out}}}n^{-1}_{e}(R)R^{2}\,dR}\\ &=V_{max}\frac{\int^{1}_{0}t^{4}(1-t)e^{-5t}\,dt}{\int^{1}_{0}t^{4}e^{-5t}\,dt}\simeq 0.31\;V_{max}\end{split}$ (12)
The mass-weighted mean outflow velocity is obtained from the maximum velocity as $v_{\mathrm{out}}=0.31V_{max}$. In addition, since the momentum rate and kinetic power are proportional to $v_{\mathrm{out}}^{2}$ and $v_{\mathrm{out}}^{3}$, respectively, we calculate the mass-weighted means of $v^{2}$ and $v^{3}$ as $\sim 0.14\;V^{2}_{max}$ and $\sim 0.07\;V^{3}_{max}$. These relations are then applied to calculate the three outflow rate parameters.
Note that the launching velocity cannot be directly measured from observational data. Instead, we test potential proxies of $V_{max}$ that can be calculated from the observed V[OIII] and $\sigma_{\mathrm{[OIII]}}$, using the mock SDSS data (Section 4.2). We find that two proxies show good agreement with the input $V_{max}$. One is a quadratic sum of $|\mathrm{V}_{\mathrm{[OIII]}}|$ and 3 $\sigma_{\mathrm{[OIII]}}$, which is similar to Bae16 or RodriguezZaurin13, and the other is a direct summation ($|\mathrm{V}_{\mathrm{[OIII]}}|$ + 3 $\sigma_{\mathrm{[OIII]}}$), which is similar to Rupke13 or Fiore17. In Figure 14, we compare the two proxies, showing that both reproduce the launching velocity with a scatter of 0.2 dex. Here, we use the first proxy (i.e., quadratic sum).
Lastly, we utilize two outflow sizes as upper and lower limits. As discussed in Section 5.3, we first adopt our $\sigma_{\mathrm{[OIII]}}$-based sizes calculated from the kinematic outflow size–luminosity relation as upper limits and determine the lower limits of outflow rates. Second, we utilize the NLR sizes calculated from the NLR size–luminosity relation, which was derived from nearby AGNs (Polack24). We determine lower limits of the outflow sizes by dividing the NLR sizes by three, which is a typical Rout/R[OIII] ratio from Fischer18; Polack24, and calculate upper limits of the outflow rates. Note that these outflow sizes are typically 0.4 - 0.6 dex smaller than the $\sigma_{\mathrm{[OIII]}}$-based sizes within the dynamic range of the [O III] luminosity.
Figure 15: Mass outflow rate (log $\dot{\mathrm{M}}_{\mathrm{out}}$, top), momentum rate (log $\dot{\mathrm{P}}_{\mathrm{out}}$, middle), and kinetic power (log $\dot{\mathrm{E}}_{\mathrm{out}}$, bottom) as a function of bolometric luminosity of the SDSS type 2 AGNs. The red and blue contours show the energetics based on the upper limits (red contours) and lower limits (blue contours) of the outflow sizes. The energetics of individual AGNs from the literature are denoted with various symbols.
In Figure 15, we compare the upper and lower limits of the outflow rates with luminosity to investigate whether the outflows are energetic enough for efficient feedback. Despite the large scatter, the estimated outflow rates (contours) fall within the ranges of $\dot{\mathrm{M}}_{\mathrm{out}}\sim 1-100\;\mathrm{L}_{\mathrm{bol}}/\eta c^{2}\;(\eta=0.1)$, $\dot{\mathrm{P}}_{\mathrm{out}}\sim 0.01-1\;\mathrm{L}_{\mathrm{bol}}/c$, and $\dot{\mathrm{E}}_{\mathrm{out}}\sim 10^{-6}-10^{-3}\;\mathrm{L}_{\mathrm{bol}}$. These ranges are consistent with previous studies, except for some discrepancies due to different definitions of outflow parameters. For example, Fiore17 sample, which has kinetic powers mostly in the range of 0.01 - 1% of Lbol, has much higher $v_{\mathrm{out}}$ values since it was defined based on the kinematics of the broad component ($|v_{\mathrm{broad}}|+2\sigma_{\mathrm{broad}}$). In contrast, we compute $v_{\mathrm{out}}$ using the first and second moments of the total [O III] profile, leading to smaller $v_{\mathrm{out}}$ at a given luminosity. This difference becomes more evident in $\dot{\mathrm{E}}_{\mathrm{out}}$ since it is proportional to $v^{3}_{\mathrm{out}}$.
Theoretical studies have suggested thresholds of kinetic power relative to bolometric luminosity (i.e., $\dot{\mathrm{E}}_{\mathrm{out}}/\mathrm{L}_{\mathrm{bol}}\sim 0.5-5\%$, e.g., Hopkins10a; Zubovas12; King15), that can suppress star formation. The estimated kinetic power is mostly much lower than the thresholds, implying that feedback may not be efficient. However, it is important to note that these outflow rates are based only on ionized outflows, which contribute a relatively low fraction of the total outflow mass (e.g., Fluetsch19). Nevertheless, 0.85% or 0.14% of the sample has $\dot{\mathrm{E}}_{\mathrm{out}}/\mathrm{L}_{\mathrm{bol}}$ higher than 0.5% when using outflow sizes based on Polack24 or Kim23 relations, respectively. These AGNs with extremely high kinetic powers also exhibit high outflow velocities, making them excellent targets for spatially resolved outflow studies in detail.
5.5 Limitations of the current model
Despite the modifications, there are still limitations in the model and the mock data analysis. First, since we adopt analytic functions for the flux and velocity distributions of outflows, our model cannot describe the detailed structures in outflows. For example, various studies of nearby AGNs found clumpy regions within the bicone structure (e.g., Venturi18; Mingozzi19; Kakkad23).
Second, our model simulates no flux errors in the mock data. For example, the $\sigma_{\mathrm{[OIII]}}$-based method adopts the radius of the furthest [O III]-detected pixel as an outflow size when the $\sigma_{\mathrm{[OIII]}}/\sigma_{*}$ does not converge to unity. In this case, the S/N becomes critical for determining the outflow size. Also, we did not test spectroastrometry-based methods (e.g., Carniani15; Singha22), which define the outflow size as a spatial offset between the center and flux centroid of the [O III] emission. Reasonable errors are necessary to fit the 2D flux map when estimating the flux centroid.
Third, we assumed that all PDFs in our MC simulations are independent since we do not have a well-described dependence among them. However, since some of the input parameters are likely to be correlated, the PDFs should be described as multivariate. Previous studies investigated the correlations between outflow parameters and bolometric luminosity (or Eddington ratio). The opening angle and outflow velocity are also suggested to have positive (Bae16), negative (Muller-Sanchez11), or no correlations (Fischer13). Therefore, MC simulations with a proper treatment of correlations between outflow parameters may better reproduce the observed gas kinematics.
6 Summary
In this study, we present the modified outflow model and its applications to local type 2 AGNs. Here we summarize the main results of this study.
  1. 1.
    We find that including a rotating disk component in the model reduces the gas velocity and velocity dispersion within the SDSS fiber size. Thus, the outflow launching velocity needs to be faster to compensate for the disk contribution. The observed VVD distribution of the local type 2 AGNs can be reproduced by $V_{max}\lesssim 1500$ km s-1.
  2. 2.
    Based on the MC simulations to reproduce the observed VVD distribution, we find that the launching velocity is $\lesssim$1000 km s-1 for a majority of the type 2 AGNs except for several percent of strong outflow cases (>> 1000 - 1500 km s-1), and that the outer half opening angle is less than 50.
  3. 3.
    By generating mock IFU data, we find similar features reported in previous IFU studies, e.g., circular morphology in the flux map and the initial plateau of the radial profiles of velocity widths. These results suggest that the biconical model with a seeing convolution can successfully reproduce observational data.
  4. 4.
    By investigating the seeing effect on the outflow size measurement, we confirmed that the outflow size measured based on velocity widths can be overestimated, particularly when the angular size of the outflow is comparable to or smaller than the seeing size. Nevertheless, this still supports the lack of global feedback since the previously reported sizes can be considered as upper limits.
This work has been supported by the Basic Science Research Program through the National Research Foundation of Korean Government (2021R1A2C3008486).

### v2

@@
 We present a modified outflow model and its application to constrain ionized outflow properties of active galactic nuclei (AGNs).
-By adding a rotating disk component to the biconical outflow model of Bae16, we find that models with a rotating disk require faster launching velocities ($\lesssim$ 1500 km s-1) than outflow-only models to be consistent with the observed gas kinematics of local type 2 AGNs.
+By adding a rotating disk component to the biconical outflow model of Bae and Woo (2016), we find that models with a rotating disk require faster launching velocities ($\lesssim$ 1500 km s-1) than outflow-only models to be consistent with the observed gas kinematics of local type 2 AGNs.
 We perform Monte Carlo simulations to reproduce the observed distribution of gas kinematics of a large sample ($\sim$ 39,000), constraining the launching velocity and opening angle.
@@
 Active galactic nuclei (16), Galaxy kinematics (602), Galaxy winds (626)
-software: astropy (Astropy22), matplotlib (matplotlib), numpy (numpy), scipy (scipy)
+software: astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020)
 show]kcs1996kcs@snu.ac.kr
-1 Introduction
-Accreting supermassive black holes (SMBHs), or active galactic nuclei (AGNs), emit a vast amount of energy through various channels, i.e., radiation, gas outflows, and relativistic jets, which are thought to interact with the interstellar medium (ISM) and eventually affect star formation in host galaxies (e.g., Silk98; Fabian12).
-Recent theoretical studies commonly implement AGN feedback to reproduce primary observables, i.e., galaxy luminosity function (e.g., Hopkins08; Weinberger17; Wellons23).
-Nonetheless, the overall role and detailed mechanism of AGN feedback remains unclear in observations (see reviews by Harrison17; Harrison24).
-AGN-driven gas outflows are widely considered as a primary candidate for a feedback channel.
-It is now well-known that these outflows are multiphase (e.g., Fiore17; Fluetsch19; Riffel23), ubiquitous (e.g., Mullaney13; Woo16; Rakshit18), and kpc-scales (e.g., Harrison14; Kim23).
-In particular, the broad wing component of the [O III] $\lambda 5007$ emission line is frequently utilized to investigate kpc-scale warm ionized gas outflows (Boroson05; Greene05).
-Large spectral surveys such as Sloan Digital Sky Survey (SDSS) enabled the scrutiny of integrated outflow properties within the central region for several tens of thousands of AGNs and their correlations with host galaxy properties (e.g., Woo17; Woo20).
+I Introduction
+Accreting supermassive black holes, or active galactic nuclei (AGNs), emit a vast amount of energy through various channels, i.e., radiation, gas outflows, and relativistic jets, which are thought to interact with the interstellar medium (ISM) and eventually affect star formation in host galaxies (e.g., Silk and Rees, 1998; Fabian, 2012).
+Recent theoretical studies commonly implement AGN feedback to reproduce primary observables, i.e., galaxy luminosity function (e.g., Hopkins et al., 2008; Weinberger et al., 2017; Wellons et al., 2023).
+Nonetheless, the overall role and detailed mechanism of AGN feedback remain unclear in observations (see reviews by Harrison, 2017; Harrison and Ramos Almeida, 2024).
+AGN-driven gas outflows are widely considered as one of the primary candidates for a feedback channel.
+It is now well-known that these outflows are multiphase (e.g., Fiore et al., 2017; Fluetsch et al., 2019; Riffel et al., 2023), ubiquitous (e.g., Mullaney et al., 2013; Woo et al., 2016; Rakshit and Woo, 2018), and kpc-scales (e.g., Harrison et al., 2014; Kim et al., 2023).
+In particular, the broad wing component of the [O III] $\lambda 5007$ emission line is frequently utilized to investigate kpc-scale warm ionized gas outflows (Boroson, 2005; Greene and Ho, 2005).
+Large spectral surveys such as Sloan Digital Sky Survey (SDSS) enabled the scrutiny of integrated outflow properties within the central region for several tens of thousands of AGNs and their correlations with host galaxy properties (e.g., Woo et al., 2017, 2020).
 On the other hand, using spatially resolved spectroscopy, many other studies have focused on the detailed outflow structures for a limited number of targets ($\lesssim$ several tens).
-One of the primary findings in early studies is that the ionized outflows are well-described by a biconical model with two axisymmetric cones (Crenshaw00a; Crenshaw00b; Veilleux01; Das05; Das06).
-Subsequent studies took into account the dust extinction by the galactic plane (e.g, Crenshaw10a; Crenshaw10b; Fischer10; Fischer11; Fischer13; Bae16) or a rotating disk (e.g., Muller-Sanchez11) to reproduce the [O III] line profile and diverse narrow line region (NLR) kinematics.
-Marconcini23 recently developed a new modeling tool for analyzing 3D gas kinematics based on the 3D distribution of numerous gas clouds.
+One of the primary findings in early studies is that the ionized outflows are well-described by a biconical model with two axisymmetric cones (Crenshaw and Kraemer, 2000; Crenshaw et al., 2000; Veilleux et al., 2001; Das et al., 2005, 2006).
+Subsequent studies took into account the dust extinction by the galactic plane (e.g, Crenshaw et al., 2010a, b; Fischer et al., 2010, 2011, 2013; Bae and Woo, 2016) or a rotating disk (e.g., Müller-Sánchez et al., 2011) to reproduce the [O III] line profile and diverse narrow line region (NLR) kinematics.
+Marconcini et al.
+(2023) recently developed a new modeling tool for analyzing 3D gas kinematics based on the 3D distribution of numerous gas clouds.
 By using a sample of nearby AGNs, these studies have provided various constraints on the geometry, extent, and energetics of outflows, which are crucial for quantifying outflow energetics and their impact on the host galaxy evolution.
 Despite the importance of estimating the outflow parameters, it is challenging to expand these works to a large sample due to observational limitations.
-For instance, spatially resolved spectroscopic observations are still expensive to be applied to a large sample.
-Furthermore, while integral field unit (IFU) surveys such as Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) or Sydney- AAO (Australian Astronomical Observatory) Multi-object Integral-field spectroscopy (SAMI) provide spatially resolved kinematics for $\lesssim$ thousands of AGNs (Wylezalek20; Deconto-Machado22; Gatto24; Oh24), they typically have a seeing-limited resolution (FWHM $\sim$ 1″- 2″), which is often insufficient to resolve outflow structures at moderate redshifts.
+For instance, spatially resolved spectroscopic observations are still too expensive to be applied to a large sample.
+Furthermore, while integral field unit (IFU) surveys such as MaNGA or SAMI provide spatially resolved kinematics for $\lesssim$ thousands of AGNs (Wylezalek et al., 2020; Deconto-Machado et al., 2022; Gatto et al., 2024; Oh et al., 2024), they typically have a seeing-limited resolution (FWHM $\sim$ 1″- 2″), which is often insufficient to resolve outflow structures at moderate redshifts.
 For example, a few kpc scales of outflows at z $\sim$ 0.1 (1 kpc $\sim$ 0.5″) are barely resolved even under the best seeing conditions on the ground.
-While space telescopes such as Hubble Space Telescope (HST) and James Webb Space Telescope (JWST) or adaptive optics (AO) can improve the spatial resolution down to a diffraction-limit (FWHM $\sim$ 0.1″), the detailed geometry and kinematics can only be resolved in rare cases such as nearby AGNs (e.g., Fischer13; Polack24) or galactic scale outflows (e.g., Vayner21a; Wylezalek22; Cresci23).
+While space telescopes such as Hubble Space Telescope (HST) and James Webb Space Telescope (JWST) or adaptive optics (AO) can improve the spatial resolution down to a diffraction-limit (FWHM $\sim$ 0.1″), the detailed geometry and kinematics can only be resolved in rare cases such as nearby AGNs (e.g., Fischer et al., 2013; Polack et al., 2024) or galactic scale outflows (e.g., Vayner et al., 2021; Wylezalek et al., 2022; Cresci et al., 2023).
 Nevertheless, detailed outflow models and simulated mock data can be efficiently utilized to overcome the limited spatial resolution and constrain the physical parameters of the outflows, which are not directly measurable.
-Bae16 constructed a 3D biconical outflow model combined with a simple extinction by a thin dust plane and investigated the effect of each parameter on the gas kinematics.
-Their main findings are as follows: (1) velocity dispersion is higher for larger intrinsic outflow velocity or bicone inclination; (2) velocity offset increases with higher dust extinction; (3) a maximum velocity of 500 - 1000 km s-1can explain most of the observed gas kinematics of local SDSS type 2 AGNs from Woo16; and (4) higher intrinsic velocities tend to have wider bicone opening angles.
+Bae and Woo (2016) constructed a 3D biconical outflow model combined with a simple extinction by a thin dust plane and investigated the effect of each parameter on the gas kinematics.
+Their main findings are as follows: (1) velocity dispersion is higher for larger intrinsic outflow velocity or bicone inclination; (2) velocity offset with respect to systemic velocity increases with higher dust extinction; (3) a maximum velocity of 500 - 1000 km s-1 can explain most of the observed gas kinematics of local SDSS type 2 AGNs from Woo et al.
+(2016); and (4) higher intrinsic velocities tend to have wider bicone opening angles.
 However, this model did not account for rotating disk emission and seeing convolution, which are critical for reproducing 2D gas kinematics.
 […]
-2 Modified 3D Biconical Outflow Model
-We summarize the original outflow model from Bae16 in Section 2.1 and demonstrate three major modifications in Section 2.2, 2.3, and 2.4, respectively.
-2.1 3D Biconical Outflow Model
-Bae16 presented a 3D outflow model adopted from Crenshaw10a; Crenshaw10b, which consists of two axisymmetric outflow cones and a simple extinction factor by a thin galactic dust plane (see Figure 1).
+II Modified 3D Biconical Outflow Model
+We summarize the original outflow model from Bae and Woo (2016) in Section II.1 and demonstrate three major modifications in Section II.2, II.3, and II.4, respectively.
+II.1 3D Biconical Outflow Model
+Bae and Woo (2016) presented a 3D outflow model adopted from Crenshaw et al.
+(2010a, b), which consists of two axisymmetric outflow cones and a simple extinction factor by a thin galactic dust plane (see Figure 1).
 Once a set of 11 parameters is given, the model constructs 3D distributions of flux and velocity.
 By collapsing the 3D model along the line-of-sight (LOS), 2D distributions of flux, velocity, and velocity dispersion are generated.
-Bae16 also calculated flux-weighted velocity and velocity dispersion using the 2D maps for a direct comparison with the observed [O III] velocity and velocity dispersion (VVD) distribution of SDSS type 2 AGNs from Woo16.
+Bae and Woo (2016) also calculated flux-weighted velocity and velocity dispersion using the 2D maps for a direct comparison with the observed [O III] velocity and velocity dispersion (VVD) distribution of SDSS type 2 AGNs from Woo et al.
+(2016).
 The input parameters define the bicone structure, 3D distributions of flux and kinematics, and the orientations of the bicone and the dust plane.
@@
 As illustrated by the dark gray region in Figure 1, we model gas outflows using a hollow bicone.
-For simplicity, we set all models in this study to have an outer half opening angle that is 20 larger than the inner one ($\theta_{out}=\theta_{in}+20^{\circ}$).
-The flux distribution is determined by e-folding radius ($\tau$) and dust extinction factor ($A$).
+For simplicity, we fix the difference between the outer and inner half opening angles to 20 ($\theta_{out}=\theta_{in}+20^{\circ}$).
+While the opening angle differences estimated by Fischer et al.
+(2013) are mostly in the 5–20 range, they adopted the minimum opening angle difference required to match the observations, meaning that the larger difference is also plausible.
+Thus, we adopted 20 as a representative value from observations.
+We also note that the inner opening angle has an opposite, but less significant effect on the integrated kinematics compared to the outer opening angle (Bae and Woo, 2016).
+The flux distribution is determined by $e$-folding radius ($\tau$), i.e., a radius where the flux decreases by a factor of $e$, and dust extinction factor ($A$).
 Since we only focus on relative flux distributions, we adopt an arbitrary central flux value ($f_{0}$).
@@
 Thus, the intrinsic 3D flux distribution can be written as $f(R)=f_{0}e^{-R/\tau}$.
-Following Bae16, we assume the e-folding radius is 20% of the outflow size ($\tau=R_{out}/5$), which gives approximately two orders of magnitude fainter flux at the outflow edge than the center.
+Following Bae and Woo (2016), we assume the $e$-folding radius is 20% of the outflow size ($\tau=R_{out}/5$), which gives approximately two orders of magnitude fainter flux at the outflow edge than at the center.
 The dust extinction factor is multiplied when collapsing the 3D flux distributions to 2D flux maps.
@@
 In all three cases, the velocity is directed radially outward from the center.
-Following Bae16, we adopt the linear decrease case for all models in this study.
+Following Bae and Woo (2016), we adopt the linear decrease case for all models in this study.
 Note that the maximum velocity can be regarded as a launching velocity, as it represents the intrinsic outflow velocity at the center.
@@
 The LOS is defined as the z-axis in this coordinate system.
-Lastly, the orientations of the model are determined by position angles (PAcone and PAdisk) and inclinations ($i_{cone}$ and $i_{disk}$) of the bicone and the dust plane (or disk in Section 2.2).
+Lastly, the orientations of the model are determined by position angles (PAs; PAcone and PAdisk) and inclinations ($i_{cone}$ and $i_{disk}$) of the bicone and the dust plane (or disk in Section II.2).
 The orientations of the bicone and the dust plane are independent of each other, allowing them to rotate freely on the sky plane.
 However, for simplicity, we fix the PAs of the bicone and dust plane to 0 and 90, respectively, meaning that the bicone axis and the major axis of the plane extend along the north-south and east-west directions, respectively.
-This is because our mock data analysis in Section 3 focuses on 3″-extracted spectra and radial profiles, which are little affected by PAs on the sky plane.
-Using this biconical model, Bae16 tested the impact of each parameter on the velocity offset and dispersion, and the line profile of the [O III] emission line.
+Since our mock data analysis focuses on 3″-extracted kinematics and radial profiles of velocity widths, the effect of PAs is negligible.
+Note that we do not consider various relative PA values between the disk and bicone for simplicity.
+We expect that the relative PA has no significant effect on mock SDSS data since we integrate all pixels within the 3″-aperture.
+In addition, velocity widths in mock IFU analysis are dominated by the bicone component in cases of strong outflows, thereby reducing the impact of the rotational orientation of the disk on the kinematics.
+The effect of relative PAs on mock IFU data will be investigated in future work.
+Using this biconical model, Bae and Woo (2016) tested the impact of each parameter on the velocity offset and dispersion, and the line profile of the [O III] emission line.
 Nevertheless, the model lacked several components crucial for comparison with observational data.
@@
 Furthermore, to accurately compare the spatial properties of the simulated outflows with observations, the seeing convolution should be taken into account.
-Therefore, we add a rotating gas component (Section 2.2) and the seeing effect (Section 2.3) to the Bae16 model.
-In addition, to improve mock data compatibility, we switch from a coordinate system normalized by the outflow size to a physical scale in arcsec units based on the redshift and outflow size parameters (Section 2.4).
-Figure 1: A schematic diagram of a biconical model example of Bae16.
-SMBH is located at the center of this plot, and the dust plane is shown as a red line between two cones.
-The dark gray area represents where the outflowing gas exists (i.e., between inner and outer opening angles), while the light gray area shows a hollow region.
-The LOS direction is shown as a black arrow on the right.
-2.2 Rotating Component
-Although the outflow is a primary kinematic component in many AGNs, the gas kinematics in the NLR are more complicated due to the presence of other effects.
-Among them, gas following the gravitational potential of the host galaxy is often observed as a narrow emission line that has similar kinematics to stellar absorption lines ($\sigma_{[\mathrm{OIII}],narrow}\sim\sigma_{*}$ Nelson96; Greene05).
-For example, if there is no non-gravitational effect, i.e., outflows, the gas kinematics are dominated by a rotating disk component (e.g., Luo19).
-Given that a majority of AGNs in the VVD diagram (see Figure 4(a)) are located at relatively small [O III] velocity offset and dispersion, the disk component should be included to properly reproduce the observed distribution of the gas kinematics.
-Even for AGNs with strong outflows, a rotational pattern is often observed in IFU data (Kang18; Kim23).
-Thus, rotation due to the host galaxy gravity is required to properly reproduce the demography of the gas kinematics as well as the characteristics in the spatially resolved kinematics.
-To generate disk models, we make use of the GalPak3D package (Galaxy Parameters and Kinematics, Bouche15), which was originally developed to model the morphokinematics of 3D observational data.
-Since our goal is to investigate how the rotating gas component influences the total gas kinematics, we employ only the basic models generated by the package without adjusting details.
-Consequently, all disk models in this study are based on the default models of GalPak3D.
-The radial and perpendicular flux profiles are set to exponential ($n=1$) and Gaussian, respectively, and the rotation curve follows a hyperbolic tan (tanh) profile.
-We regulate two flux parameters and three kinematic parameters of the disk model.
+Therefore, we add a rotating gas component (Section II.2) and the seeing effect (Section II.3) to the Bae and Woo (2016) model.
+In addition, to improve mock data compatibility, we switch from a coordinate system normalized by the outflow size to a physical scale in arcsec units based on the redshift and outflow size parameters (Section II.4).
+Figure 1: Example biconical outflow model from Bae and Woo (2016).
+The dust plane is shown as a red line between the two cones.
+The dark gray region represents the volume occupied by the outflowing gas (i.e., between inner and outer opening angles), while the light gray region denotes the hollow interior.
+The LOS direction is indicated by the black arrow.
+The definitions and adopted values of the cone and disk inclination angles are also shown.
+II.2 Rotating Component
+While outflows are the dominant nongravitational component in AGN kinematics, the gas motions in the NLR are more complex due to additional dynamical effects.
+In particular, gas motion due to the gravitational potential of the host galaxy is represented by a core component of narrow emission lines, which show similar line width and offset compared to stellar absorption lines ($\sigma_{[\mathrm{OIII}],narrow}\sim\sigma_{*}$ Nelson and Whittle, 1996; Greene and Ho, 2005).
+If there is no nongravitational effect, gas kinematics are governed by a host galaxy potential, i.e., a rotating disk (e.g., Luo et al., 2019).
+Given that a majority of AGNs in the VVD diagram (see Figure 4(a)) show relatively small velocity offset and dispersion, the disk component should be included to reproduce the observed distribution of the gas kinematics properly.
+Even for AGNs with strong outflows, a rotational pattern is often observed in IFU data (Kang and Woo, 2018; Kim et al., 2023).
+Thus, a disk component is required to reproduce the characteristics in the spatially resolved kinematics.
+To generate disk models, we use the GalPak3D package (Galaxy Parameters and Kinematics, Bouché et al., 2015), which was originally developed for the morphokinematics of 3D observational data.
+Since our goal is to investigate the effect of a rotating disk on the observed gas kinematics, we employ only the basic models generated by the package without adjusting details.
+Using the default models of GalPak3D, we assume radial and perpendicular flux profiles as exponential ($n=1$) and Gaussian, respectively, and set the rotation curve to a hyperbolic tan (tanh) profile.
+We regulate two flux parameters ($f_{disk}/f_{cone}$ and $r_{1/2}$) and three kinematic parameters ($V_{max,rot}$, $r_{t}$, and $\sigma_{0}$) of the disk model.
 The inclination and PA of the disk are assumed to be the same as those of the dust plane in the outflow model.
@@
 $v(r)=V_{max,rot}\;\mathrm{tanh}(r/r_{t}),$ (3)
-where $r$ is the radius on the galactic plane. In addition, we adjust the intrinsic velocity dispersion ($\sigma_{0}$), which corresponds to the dynamical hotness of the disk.
+where $r$ is the radius on the galactic plane.
+At $r>r_{t}$, the rotational velocity begins to follow a flat radial profile.
+In addition, we adjust the intrinsic velocity dispersion ($\sigma_{0}$), which corresponds to the dynamical hotness of the disk.
 Once the five disk parameters and the disk orientation (PAdisk and $i_{disk}$) are given, GalPak3D computes flux, velocity, and velocity dispersion maps.
@@
 These combined maps demonstrate that the biconical structure is properly embedded in the background disk model.
-2.3 Seeing Effect
-The beam smearing effect by atmospheric seeing has been pointed out as a significant obstacle in spatially resolved studies of AGNs (e.g., Husemann16; Villar-Martin16; Singha22).
+II.3 Seeing Effect
+The beam smearing effect by atmospheric seeing has been pointed out as a significant obstacle in spatially resolved studies of AGNs (e.g., Husemann et al., 2016; Villar-Martín et al., 2016; Singha et al., 2022).
 Given that the seeing effect is more pronounced at higher redshifts, we need to include it in our model to account for outflows in a majority of AGN host galaxies.
@@
 The seeing convolution blurs the central bicone structure, so the final images are more like a circular shape in all 2D maps.
-We will discuss the seeing effect in Section 4.3 and 5.3, focusing on implications for the outflow size measurements in IFU data.
-2.4 Mock Data Compatibility
-We update our model to calculate 2D maps in an (x, y) coordinate system defined in arcsec units based on the input redshift and physical size of the outflow, while Bae16 model adopted relative scales normalized by the outflow size.
+We will discuss the seeing effect in Section IV.3 and V.3, focusing on implications for the outflow size measurements in IFU data.
+II.4 Mock Data Compatibility
+We update our model to calculate 2D maps in an (x, y) coordinate system defined in arcsec units based on the input redshift and physical size of the outflow, while the Bae and Woo (2016) model adopted relative scales normalized by the outflow size.
 This approach offers several advantages: (1) the outflow model uses the same pixel units as the disk model, making it convenient to combine the two models; (2) mock data can be directly compared with real observational data.
 Accordingly, we specify the mock data by pixel size and field-of-view (FoV) parameters, and all calculations are performed using this pixel scale from given redshift and physical length parameters ($R_{out}$, $r_{1/2}$, and $r_{t}$).
-Figure 2: 2D map examples of the modified model.
-From left to right, three 2D maps in each row present flux, velocity, and velocity dispersion of the corresponding model.
-First row: Bae16 biconical model.
-This model has the same geometry as Figure 1.
-Second row: Disk model from GalPak3D.
-Third row: The bicone + disk combined model.
-Last row: The combined model after seeing convolution.
-Red dashed circles at the center and red bars on the top right show the FWHM seeing size.
-We also plot 3″-aperture size as gray circles, which is the fiber size of SDSS spectra.
-3 Mock Data Analysis
+Figure 2: Examples of 2D maps produced by the modified model.
+From left to right, each column shows the flux, velocity, and velocity dispersion maps.
+The first row presents the biconical outflow model of Bae and Woo (2016).
+The second row shows the rotating disk model generated with GalPak3D.
+The third row displays the combined bicone + disk model, and the last row presents the combined model after applying the seeing convolution.
+Red dashed circles at the center, and red bars in the upper-right corner indicate the FWHM of the seeing.
+Gray circles denote the 3″ aperture corresponding to the SDSS fiber size.
+III Mock Data Analysis
 In this Section, we describe our main analysis procedures based on mock data generated from the modified outflow model.
-First, we generate mock SDSS data extracted from the central 3″-aperture, and compare them with the observed [O III] kinematics of a large sample of SDSS type 2 AGNs in Section 3.1.
-In Section 3.2, we produce mock IFU data to analyze outflow sizes and 2D distributions of flux, velocity, and velocity dispersion.
-3.1 Mock SDSS Data
-Bae16 integrated the [O III] velocity offset and dispersion within the outflow radius and compared them with the observed VVD distribution of $\sim$39,000 SDSS type 2 AGNs from Woo16.
+First, we generate mock SDSS data extracted from the central 3″-aperture, and compare them with the observed [O III] kinematics of a large sample of SDSS type 2 AGNs in Section III.1.
+In Section III.2, we produce mock IFU data to analyze outflow sizes and 2D distributions of flux, velocity, and velocity dispersion.
+III.1 Mock SDSS Data
+Bae and Woo (2016) integrated the [O III] velocity offset and dispersion within the outflow radius and compared them with the observed VVD distribution of $\sim$39,000 SDSS type 2 AGNs from Woo et al.
+(2016).
 Here, we use our modified model to reexamine the comparison with the observed VVD distribution.
@@
 Thus, we mainly investigate the effect of the rotating gas component on the VVD distribution.
-3.1.1 Model Grids for the Observed VVD
+III.1.1 Model Grids for the Observed VVD
 To investigate the effect of outflow properties on the observed kinematics, we overlay various model grids on the observed VVD diagram.
@@
 For each model grid, we also adjust the dust extinction rate from 0 to 100 % (i.e., $A$ = 1 to 0) and launching velocities by 500, 1000, and 2000 km s-1.
-3.1.2 Mock VVD from Monte Carlo Simulation
+III.1.2 Mock VVD from Monte Carlo Simulation
 We perform Monte Carlo (MC) simulations and generate mock VVD distributions by assuming a probability distribution function (PDF) of each parameter.
@@
 Specifically, when $|\sigma_{\mathrm{[OIII]}}-\sigma_{*}|$ is less than 30 km s-1, we classify this single Gaussian as disk-only emission, i.e., $f_{disk}/f_{cone}=\infty$, otherwise as outflow-only emission, i.e., $f_{disk}/f_{cone}=0$.
-We also estimate the PDF of the dust extinction factor by using observed $\alpha/H$$\beta$ with $R^{A}_{V}=4.5$ (Vogt13).
+We also estimate the PDF of the dust extinction factor by using observed $\alpha/H$$\beta$ with $R^{A}_{V}=4.5$ (Vogt et al., 2013).
 Note that we restrict the dust extinction rate to smaller than 80% (or $A>0.2$) since a higher extinction rate results in too broad mock VVD distributions along the V[OIII] axis, which are inconsistent with the observed distribution.
@@
 The stellar mass is taken from the MPA-JHU catalog  [^1] https://wwwmpa.mpa-garching.mpg.de/SDSS/ and divided into five bins ($\rm{log}\;M_{*}=$9.7 to 11.3 with 0.4 step).
-We calculate the four remaining disk parameters, i.e., $r_{1/2}$ (Dutton11), $V_{max,rot}$ (Torres-Flores11), $\sigma_{0}$ (Oh22), and $r_{t}$ (Amorisco10), using empirical relations with stellar mass.
+We calculate the four remaining disk parameters, i.e., $r_{1/2}$ (Dutton et al., 2011), $V_{max,rot}$ (Torres-Flores et al., 2011), $\sigma_{0}$ (Oh et al., 2022), and $r_{t}$ (Amorisco and Bertin, 2010), using empirical relations with stellar mass.
 Thus, once the stellar mass is selected among the five values, these four disk parameters are determined accordingly.
@@
 For the outflow model, we first adopt random inclinations and fixed PAs (PAcone=0 and PAdisk=90) for both the bicone and dust plane.
-The inclination of the dust plane is given from 7.5 to 172.5 with 15 step, and its PDF is proportional to the solid angle of the normal vector integrated from 0 to 360 of PA.
-Since we only deal with type 2 AGNs, the bicone inclination is restricted to less than 40, i.e., 7.5, 22.5, and 37.5, and its PDF is also proportional to the solid angle of the bicone axis.
+The inclination of the dust plane is given from 7.5 to 172.5 with a 15 step, and its PDF is proportional to the solid angle of the normal vector integrated from 0 to 360 of PA.
+Since we focus on type 2 AGNs, we restrict the bicone inclination to less than 40 (7.5, 22.5, and 37.5), and its PDF is also proportional to the solid angle of the bicone axis.
+Models with higher bicone inclinations, i.e., more aligned bicone axis with the LOS, are likely to be observed as type 1 AGNs, as their accretion disk and broad line region (BLR) are readily exposed to the observer (e.g., Marin, 2014).
 Then we assume that the PDF for five outflow size values (0.3, 1, 3, 6, 9 kpc) follows a Gaussian distribution with a mean of zero and a standard deviation of 2 kpc.
@@
 For these two parameters, we assume that their PDFs follow a truncated Gaussian function.
-We first set the range of the launching velocity to $100,\;\infty)$ km s-1  and restrict the outer opening angle values from 30 to 70 with a 10 step.
+We first set the range of the launching velocity to $100,\;\infty)$ km s-1 and restrict the outer opening angle values from 30 to 70 with a 10 step based on Fischer et al.
+(2013).
+Note that larger outer opening angles are excluded due to their physically extreme geometry, while smaller opening angles are excluded since they do not improve the main results.
 We then run multiple MC simulations by changing the mean and standard deviation of the Gaussian to find the best $V_{max}$ and $\theta_{out}$ distributions that generate the best-matching mock VVD with the observed VVD.
 Each MC simulation generates 100,000 pairs of velocity offset and dispersion.
-3.1.3 Kullback-Leibler Divergence and $D_{max}$
+III.1.3 Kullback-Leibler Divergence and $D_{max}$
 To quantitatively compare the mock VVDs with the observed VVD, we apply two statistical quantities that assess proximity between two 2D distributions.
-We first calculate Kullback-Leibler divergence ($D_{KL}(P||Q)$, KLDIV) as follows:
+We first calculate Kullback-Leibler (KL) divergence ($D_{KL}(P||Q)$, Kullback and Leibler, 1951) as follows:
 $D_{KL}(P||Q)=\sum P(v_{obs},\sigma_{obs})\;\mathrm{log}\frac{P(v_{obs},\sigma_{obs})}{Q(v_{int},\sigma_{int})},$ (8)
 where $P(v_{obs},\sigma_{obs})$ and $Q(v_{int},\sigma_{int})$ are probability distributions of the observed and mock VVD, respectively.
-In brief, the Kullback-Leibler divergence is a measure of a statistical distance between two PDFs derived from information theory, i.e., the smaller the Kullback-Leibler divergence is, the more similar the two PDFs are.
+In brief, the KL divergence is a measure of a statistical distance between two PDFs derived from information theory, i.e., the smaller the KL divergence is, the more similar the two PDFs are.
 The second quantity is the $D_{max}$ from the 2D Kolmogorov-Smirnov test, calculated from the public code ndtest [^2] Written by Zhaozhou Li, https://github.com/syrte/ndtest.
@@
 Note that we add random error smoothing to all mock VVDs based on typical SDSS uncertainties of velocity offset and dispersion to account for the observational uncertainties.
-To remove variations induced by this error smoothing, we repeat this process 100 times in each grid of $V_{max}$ and $\theta_{out}$ PDFs and take the mean and standard deviation of Kullback-Leibler divergence and $D_{max}$ values.
-If the mean Kullback-Leibler divergence (or $D_{max}$) for given parameter distributions is notably smaller than for the others, by more than its standard deviation, then that mock VVD distribution is considered a better reproduction of the observed data.
-Figure 3: Histograms of physical properties of Woo16 type 2 AGNs.
-From left to right, each histogram shows redshift, stellar mass, flux ratio between disk and outflow components, and extinction factor based on Balmer decrement.
-3.2 Mock IFU Data
+To remove variations induced by this error smoothing, we repeat this process 100 times in each grid of $V_{max}$ and $\theta_{out}$ PDFs and take the mean and standard deviation of KL divergence and $D_{max}$ values.
+If the mean KL divergence (or $D_{max}$) for given parameter distributions is notably smaller than for the others, by more than its standard deviation, then that mock VVD distribution is considered a better reproduction of the observed data.
+Figure 3: Histograms of physical properties of the type 2 AGN sample from Woo et al.
+(2016).
+From left to right, the panels show the distributions of redshift, stellar mass, flux ratio between disk and outflow components, and extinction factor derived from the Balmer decrement.
+III.2 Mock IFU Data
 We simulate mock 2D maps using various redshifts and outflow sizes to test the validity of previous methods for outflow size measurements.
 In particular, we focus on the dependence of the measured sizes on the seeing size normalized by the angular size of the outflow.
-In our previous IFU studies, we developed a kinematic approach to measure the outflow size and applied it to AGNs at $z<0.3$ (Karouzos16a; Kang18; Bae17; Luo21; Kim23).
+In our previous IFU studies, we developed a kinematic approach to measure the outflow size and applied it to AGNs at $z<0.3$ (Karouzos et al., 2016; Kang and Woo, 2018; Bae et al., 2017; Luo et al., 2021; Kim et al., 2023).
 The outflow region is defined as where the [O III] velocity dispersion is more enhanced than stellar velocity dispersion, which traces the gravitational potential of the host galaxies.
@@
 Since we do not have stellar velocity dispersion in our model, we alternatively adopt the [O III] velocity dispersion extracted within a 3″-aperture of the convolved disk-only maps.
-This velocity dispersion of the disk model mimics the [O III] narrow/core component, which is an optimal proxy of stellar velocity dispersion (e.g., Greene05).
-Another widely used method for outflow size measurements is W80-based method (e.g., Harrison14; Wylezalek20; Ruschel-Dutra21), where W80 is defined as a [O III] line width including 80% of the total line flux.
+This velocity dispersion of the disk model mimics the [O III] narrow/core component, which is an optimal proxy of stellar velocity dispersion (e.g., Greene and Ho, 2005).
+Another widely used method for outflow size measurements is W80-based method (e.g., Harrison et al., 2014; Wylezalek et al., 2020; Ruschel-Dutra et al., 2021), where W80 is defined as a [O III] line width including 80% of the total line flux.
 Similar to the $\sigma_{\mathrm{[OIII]}}$-based method, the W80-based method defines the outflow region where the W80 is broader than a certain criterion, frequently set to 600 km s-1.
@@
 […]
-4 Result
-4.1 Location of Model Grids in VVD Diagram
-Figure 4: Model grids from the modified outflow model.
-Each color indicates a different launching velocity, 500 km s-1(cyan), 1000 km s-1(yellow), and 2000 km s-1(red): (a) The observed [O III] VVD distribution of type 2 AGNs (Woo16).
-This color map is also plotted as contours and gray color maps in the other three panels; (b) Model grids with (solid lines) and without (dashed lines) the disk.
-Each line shows the change in kinematics depending on extinction at a fixed inclination; (c) Model grids showing the effects of bicone inclinations and extinction rates when the disk component is included.
-Solid lines indicate variations with extinction at a fixed bicone inclination, while dotted lines indicate variations with inclination at a fixed extinction rate; (d) Model grids with different opening angles and extinction rates.
-Each solid and dotted line indicates the same bicone opening angle and extinction rate, respectively.
+IV Result
+IV.1 Location of Model Grids in VVD Diagram
+Figure 4: Model grids with a fixed launching velocity of 500 km s-1 (cyan), 1000 km s-1 (yellow), and 2000 km s-1 (red).
+(a) Observed [O III] VVD distribution of 39,000 type 2 AGNs from Woo et al.
+(2016); the color scale indicates the number of AGNs in each bin.
+This distribution is also shown as contours and gray heat maps in panels (b)–(d).
+(b) Model grids with (solid lines) and without (dashed lines) the disk component.
+Each line traces the change in kinematics as a function of the extinction at a fixed inclination.
+(c) Model grids showing the effects of bicone inclination and extinction when the disk component is included.
+Solid lines indicate variations with extinction at a fixed inclination, while dotted lines indicate variations with inclination at a fixed extinction.
+(d) Model grids with different opening angles and extinction values.
+For clarity, solid and dotted lines denote models with fixed opening angles and fixed extinction values, respectively.
 We investigate the dependency of outflow parameters on the location in the VVD diagram by comparing model grids with observations.
-The observed VVD distribution of local type 2 AGNs in Figure 4(a) is characterized by a V-shaped envelope and more AGNs with blueshifted [O III] than redshifted (Woo16).
-First, we examine the effect of the rotating disk emission by comparing model grids with/without the disk component (Figure 4(b)), finding that both the integrated velocity offset and dispersion ($v_{int}$ and $\sigma_{int}$) decrease by including a disk component.
-This is because the disk component usually has a smaller velocity offset and dispersion than the outflow, resulting in smaller velocity offset and dispersion values in the combined model than in the outflow-only model.
-Thus, the modified model with a disk component requires a higher launching velocity ($\lesssim$ 1500 km s-1) to reproduce a similar VVD range, while most of AGNs in the VVD diagram were explained by $V_{max}\lesssim 1000$ km s-1(orange dashed lines) in the original model (Bae16).
-Second, we revisit the effect of the bicone inclination and the dust extinction (Figure 4(c)).
-We find that both $\sigma_{int}$ and $v_{int}$ increase at a more inclined bicone, which is consistent with Bae16.
-For higher extinction, $\sigma_{int}$ decreases while $v_{int}$ increases because of the less contribution of the receding cone.
-Notably, the angle of the V-shape becomes smaller than that of the outflow-only model since the ratio of $|v_{int}|/\sigma_{int}$ decreases due to the zero velocity offset of the disk component.
+The observed VVD distribution of local type 2 AGNs (Figure 4(a)) is characterized by a V-shaped envelope and a larger number of blueshifted than redshifted [O III] sources (Woo et al., 2016).
+First, we examine the influence of the rotating disk component by comparing model grids with and without disk emission (Figure 4(b)).
+Including the disk component reduces both the integrated velocity offset and dispersion ($v_{\rm int}$ and $\sigma_{\rm int}$), because the disk typically exhibits smaller velocity offsets and dispersions than the outflow.
+As a result, the combined model requires a higher launching velocity up to 1500 km s-1 to reproduce the observed VVD distribution, whereas the original outflow-only model (Bae and Woo, 2016) could explain most AGNs with $V_{\rm max}$ up to 1000 km s-1 (orange dashed lines).
+Second, we revisit the effects of bicone inclination and dust extinction (Figure 4(c)).
+Both $v_{\rm int}$ and $\sigma_{\rm int}$ increase with larger inclination angles, consistent with Bae and Woo (2016).
+Increasing extinction decreases $\sigma_{\rm int}$ but increases $v_{\rm int}$, owing to the reduced contribution from the receding cone.
+Notably, the $|v_{\rm int}|/\sigma_{\rm int}$ ratio is generally smaller than in the outflow-only model because the disk component has zero intrinsic velocity offset.
 Lastly, we investigate the effect of the opening angle (Figure 4(d)).
-We find that a wider opening angle generates a larger $\sigma_{int}$ due to a broader range of LOS velocities, while $v_{int}$ is almost constant regardless of the opening angle.
-Overall, the effect of the opening angle is consistent with Bae16.
-4.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
-We constrain the distributions of $V_{max}$ and $\theta_{out}$ parameters by comparing mock VVD distributions from the MC simulations with the observed VVD distribution.
-We conduct 315 realizations of MC simulations using Gaussian distributions, for which the mean and standard deviation of $V_{max}$ change from 150 to 450 km s-1and from 250 to 450 km s-1, respectively, in 50 km s-1steps.
-Mean and standard deviation of the $\theta_{out}$ parameter are set to 30, 40, 50 and 5, 10, 20, respectively.
-Note that the adopted mean and standard deviation are different from those of the resulting PDFs, as we are using a truncated Gaussian distribution.
-Figure 5: Left: Kullback-Leibler divergence values depending on the PDF of the launching velocity.
-The PDF of the opening angle is fixed to the best.
-The mean and standard deviation of $V_{max}$ first determine the Gaussian distribution, and we use the truncated Gaussian from 100 km s-1as a PDF of the $V_{max}$.
-The minimum Kullback-Leibler divergence is found when the mean and standard deviation are 250 and 400 km s-1.
-Right: The same plot but $D_{max}$ from the 2D KS test.
-The minimum value is found when the mean and standard deviation are 350 and 300 km s-1.
-Figure 6: Left: Kullback-Leibler divergence values depending on the PDF of the opening angle.
-The minimum Kullback-Leibler divergence is found when the mean and standard deviation are 30 and 10, respectively.
-Right: The same plot but $D_{max}$ from the 2D KS test.
-The minimum value is found when the mean and standard deviation are 30 and 5.
-We first investigate variations of Kullback-Leibler divergence and $D_{max}$ values depending on the input launching velocity and outer opening angle PDFs in Figure 5 and 6.
-In each map, we fix the PDF of the other parameter to the best one.
-The Kullback-Leibler divergence and $D_{max}$ maps of the launching velocity show local minima near the center, ruling out the lower left and upper right corners.
-This indicates that the observed VVD distribution prefers a moderate range of launching velocity in general.
-On the other hand, the PDF of the opening angle prefers small values (i.e., lower left corner regions in Figure 6).
-Specifically, Kullback-Leibler divergence is minimized when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(250 km s-1, 400 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 10).
+A wider opening angle produces a larger $\sigma_{\rm int}$ due to the broader distribution of LOS velocities, while $v_{\rm int}$ remains nearly constant.
+Overall, the dependence on opening angle is consistent with the trends reported by Bae and Woo (2016).
+IV.2 Constraints on $V_{max}$ and $\theta_{out}$ Distributions
+We constrain the distributions of $V_{\rm max}$ and $\theta_{\rm out}$ by comparing the simulated VVD distributions with the observed VVD distribution.
+We perform multiple MC realizations, adopting Gaussian PDFs in which the mean and standard deviation of $V_{\rm max}$ range from 150 to 450 km s-1 and from 250 to 450 km s-1, respectively, in steps of 50 km s-1.
+For $\theta_{\rm out}$, we use mean values of $30^{\circ}$, $40^{\circ}$, and $50^{\circ}$, and standard deviations of $5^{\circ}$, $10^{\circ}$, and $20^{\circ}$.
+We note that the adopted means and standard deviations do not correspond exactly to those of the resulting PDFs because the distributions are truncated Gaussians.
+We then calculate 4D arrays containing 315 values ($7\times 5\times 3\times 3$) for the KL divergence and $D_{max}$, and identify the best PDFs where these values are minimized.
+As a result, KL divergence is minimized when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(250 km s-1, 400 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 10).
 $D_{max}$ exhibits a minimum value when the launching velocity follows ($\mu_{V},\sigma_{V}$)=(350 km s-1, 300 km s-1) and the opening angle follows ($\mu_{\theta},\sigma_{\theta}$)=(30, 5).
-We present the best mock VVD distributions based on Kullback-Leibler divergence (left) and $D_{max}$ (right) in Figure 7.
+Figure 5: Left: KL divergence values for differently assumed PDFs of the launching velocity $V_{\rm max}$, while keeping the opening angle distribution fixed to its best-fit value.
+The Gaussian parameters (mean and standard deviation) for $V_{\rm max}$ are varied and truncated below 100 km s-1.
+The minimum KL divergence is found at a mean of 250 km s-1 and a standard deviation of 400 km s-1.
+Right: same plot but showing $D_{\rm max}$ values instead of KL divergence.
+In this case, the minimum occurs at a mean of 350 km s-1 and a standard deviation of 300 km s-1.
+In each panel, the location of the lowest value is highlighted by a dashed-line box.
+Figure 6: Left: KL divergence values computed for differently assumed PDFs of the opening angle $\theta_{\rm out}$.
+The minimum KL divergence is found when the mean and standard deviation are $30^{\circ}$ and $10^{\circ}$, respectively.
+Right: Same as left, but showing the $D_{\rm max}$ values from the 2D KS test.
+The minimum occurs at a mean of $30^{\circ}$ and a standard deviation of $5^{\circ}$.
+In each panel, the location of the lowest value is highlighted by a dashed-line box.
+We first investigate variations in the KL divergence and $D_{max}$ values depending on the input launching velocity and outer opening angle PDFs in Figures 5 and 6.
+For visualization, we collapse the 4D arrays into 2D maps by fixing the other parameter to its best PDF.
+The KL divergence and $D_{max}$ maps of the launching velocity show local minima near the center, ruling out the lower left and upper right corners.
+This indicates that the observed VVD distribution generally prefers a moderate range of launching velocity.
+On the other hand, the opening angle PDF favors small values (i.e., lower left corner regions in Figure 6).
+We present the best mock VVD distributions based on KL divergence (left) and $D_{max}$ (right) in Figure 7.
 The two best mock VVD distributions are generally consistent with the observed one, reproducing both the V-shape feature and the overall range of the V[OIII] and $\sigma_{\mathrm{[OIII]}}$.
-The best mock VVD distribution based on Kullback-Leibler divergence exhibits a slightly larger $\sigma_{\mathrm{[OIII]}}$ range than $D_{max}$.
+The best mock VVD distribution based on KL divergence exhibits a slightly larger $\sigma_{\mathrm{[OIII]}}$ range than $D_{max}$.
 Despite the extinction limit ($A\geq 0.2$), the V-shape envelope of the mock VVD distribution is slightly broader than the observed one, especially in the low $\sigma_{\mathrm{[OIII]}}$ regime.
 This may result from several factors, such as large observational uncertainties for weak outflow AGNs.
-Figure 7: The best mock VVD distributions based on the Kullback-Leibler divergence (left) and $D_{max}$ (right).
-The contours from the observed VVD distribution are overlaid.
-Dashed box indicates a selection criterion of strong outflow AGNs ($|\mathrm{V}_{\mathrm{[OIII]}}|>300$ km s-1or $\sigma_{\mathrm{[OIII]}}>400$ km s-1).
-Figure 8: The best PDFs of the launching velocity based on the Kullback-Leibler divergence (left) and $D_{max}$ (right).
-The histograms show the best PDFs for the entire mock VVD distributions, which are derived from Gaussian distributions truncated from 100 km s-1.
-The mean and standard deviation of the Gaussian are denoted in each panel.
-Figure 9: The launching velocity distributions for a subset of strong outflow targets (outside the dashed boxes in Figure 7).
-The distributions are obtained from the best mock VVD distributions based on Kullback-Leibler divergence (left) and $D_{max}$ (right).
-Figure 10: The best PDFs of the opening angle based on the Kullback-Leibler divergence (left) and $D_{max}$ (right). Both follow Gaussian distributions truncated from 30 to 70.
-The histograms in Figure 8 show the best PDFs of the launching velocity based on Kullback-Leibler divergence (left) and $D_{max}$ (right).
+Figure 7: Best-match mock VVD distributions selected based on KL divergence (left) and $D_{\rm max}$ (right).
+The contours from the observed VVD distribution are overlaid for comparison.
+The dashed-box outlines the selection criterion for “strong outflow” AGNs, defined as $|\mathrm{V}_{\rm[OIII]}|>300\ \mathrm{km\ s^{-1}}$ or $\sigma_{\rm[OIII]}>400\ \mathrm{km\ s^{-1}}$.
+Figure 8: Best-fitting PDFs of the launching velocity based on the KL divergence (left) and $D_{max}$ (right).
+Histograms show the adopted Gaussian PDFs for the full mock VVD distributions with truncation at 100 km s-1.
+The mean and standard deviation of each Gaussian are indicated in their respective panels.
+Figure 9: Distributions of the launching velocity for a subsample of strong-outflow AGNs (those lying outside the dashed boxes in Figure 7).
+Histograms are derived from the best-matching mock VVD distributions based on KL divergence (left) and $D_{\rm max}$ (right).
+Figure 10: Best-fit PDFs for the outflow opening angle ($\theta_{\rm out}$) based on the KL divergence (left) and $D_{max}$ (right).
+All distributions are truncated Gaussians defined over the range $30^{\circ}$$70^{\circ}$.
+The histograms in Figure 8 show the best PDFs of the launching velocity based on the KL divergence (left) and $D_{max}$ (right).
 The two quantities adopt similar best PDFs for the launching velocity.
-As predicted in Section 4.1, the launching velocities are mostly smaller than 1000 km s-1, while 2 - 5% exhibit higher launching velocities (1000 - 1500 km s-1).
+As predicted in Section IV.1, the launching velocities are mostly smaller than 1000 km s-1, while 2 - 5% exhibit higher launching velocities (1000 - 1500 km s-1).
 In contrast, strong outflow AGNs exhibit significantly different $V_{max}$ distribution from the total sample.
@@
 Lastly, the best PDFs of the opening angle (Figure 10) are slightly different depending on the statistical quantities.
-Specifically, $\theta_{out}\geq 50^{\circ}$ is only exhibited in the best PDF based on Kullback-Leibler divergence, while the best PDF based on $D_{max}$ is concentrated on $\theta_{out}=30^{\circ}-40^{\circ}$.
+Specifically, $\theta_{out}\geq 50^{\circ}$ is only exhibited in the best PDF based on the KL divergence, while the best PDF based on $D_{max}$ is concentrated on $\theta_{out}=30^{\circ}-40^{\circ}$.
 Nevertheless, the two best PDFs consistently favor relatively narrow outer opening angles, indicating that the opening angle of the local type 2 AGN is likely to be narrow, mostly 30 - 40.
-4.3 Testing Outflow Sizes
+IV.3 Testing Outflow Sizes
 Table 1: Model Parameters for Mock IFU Analysis
@@
 — (1) Redshift; (2) outflow size; (3) maximum (launching) velocity; (4) outer opening angle; (5) bicone inclination; (6) disk (and dust plane) inclination; (7) extinction factor (smaller value means more extinction); (8) flux ratio between disk and bicone components; (9) velocity turnover radius (disk); (10) intrinsic velocity dispersion (disk); (11) maximum rotational velocity (disk); (12) effective radius.
-Figure 11: An example of mock IFU analysis when redshift is 0.05.
-Left: Radial profiles of $\sigma_{\mathrm{[OIII]}}$ (yellow markers) and W80 (blue markers).
-Large colored markers show the means of velocity widths in individual radial bins.
-Blue and yellow horizontal lines represent the criteria of the outflow region for W80 (600 km s-1) and $\sigma_{\mathrm{[OIII]}}$ ($\sigma_{*}$), respectively.
-Vertical dashed lines indicate half-width-half-maximum (HWHM) seeing size (red), input outflow size (black), the W80-based size (blue), and the $\sigma_{\mathrm{[OIII]}}$-based size (yellow).
-Right: 2D maps of velocity dispersion and W80.
-The colors of dashed line circles follow the same color scheme as vertical lines in the left panel.
-Figure 12: The same plots as Figure 11 but when redshift is 0.3.
+[…]
+Figure 12: Same as Figure 11 but for a higher redshift of $z=0.3$.
 We investigate the validity of the outflow size measurements based on mock IFU data by testing the effect of seeing.
@@
 We present the spatial distributions of velocity widths of the two examples in Figure 11 and 12.
-In our previous IFU studies, radial profiles of velocity dispersion often exhibit constant values in the central region (i.e., initial plateau) before a smooth decrease (e.g., Karouzos16a; Kang18).
+In our previous IFU studies, radial profiles of velocity dispersion often exhibit constant values in the central region (i.e., initial plateau) before a smooth decrease (e.g., Karouzos et al., 2016; Kang and Woo, 2018).
 This initial plateau is well reproduced in our mock IFU data, for both $\sigma_{\mathrm{[OIII]}}$ (yellow) and W80 (blue).
@@
 While the outflow size measured based on the emission line width is reliable when the seeing size is relatively small, the outflow size can be overestimated when the seeing size is similar to or larger than the angular size of the outflow.
-5 Discussion
-5.1 Comparison with Other Outflow Models
+V Discussion
+V.1 Comparison with Other Outflow Models
 In this section, we discuss the characteristics of our model in comparison with the previous models developed over the last two decades.
 First, early outflow models did not include a rotating disk.
-For example, Crenshaw00a; Crenshaw00b constructed a hollow bicone model to reproduce position-velocity (PV) diagrams of nearby AGNs (e.g., Das05; Das06).
-Subsequently, Crenshaw10a; Crenshaw10b added a dust extinction component to the model to apply for dozens of nearby AGNs (e.g., Fischer10; Fischer11; Fischer13; Meena21; Falcone24).
+For example, Crenshaw and Kraemer (2000); Crenshaw et al.
+(2000) constructed a hollow bicone model to reproduce position-velocity (PV) diagrams of nearby AGNs (e.g., Das et al., 2005, 2006).
+Subsequently, Crenshaw et al.
+(2010a, b) added a dust extinction component to the model to apply for dozens of nearby AGNs (e.g., Fischer et al., 2010, 2011, 2013; Meena et al., 2021; Falcone et al., 2024).
 Since these models were developed for the HST observations of nearby AGNs, it is not essential to include the effect of the point spread function (PSF).
-On the other hand, Carniani15 developed a model with a single approaching cone for testing their spectroastrometric analysis of powerful outflows in quasars at cosmic noon.
+On the other hand, Carniani et al.
+(2015) developed a model with a single approaching cone for testing their spectroastrometric analysis of powerful outflows in quasars at cosmic noon.
 Since the seeing effect is substantial for their high-z AGNs, they convolved each velocity channel map with the PSF, reproducing consistent [O III] flux, velocity, and velocity dispersion maps with the observations.
@@
 Second, other models were also reported for accommodating various kinematical considerations.
-For example, Muller-Sanchez11 constructed a model to analyze the coronal line regions (CLRs) using AO-assisted near-infrared IFU data.
+For example, Müller-Sánchez et al.
+(2011) constructed a model to analyze the coronal line regions (CLRs) using AO-assisted near-infrared IFU data.
 By combining a bicone model with a rotating disk, they successfully fitted CLR kinematics at the central several hundred pc, while no seeing effect was included.
-Recently, Marconcini23 developed a new model (MOKA3D) based on Venturi17 by constructing a 3D distribution of point-like clouds to reproduce the observed IFU data.
+Recently, Marconcini et al.
+(2023) developed a new model (MOKA3D) based on Venturi et al.
+(2017) by constructing a 3D distribution of point-like clouds to reproduce the observed IFU data.
 The model also convolves the 3D data and assigns a proper weight to each cloud.
-Since outflows often exhibit clumpy regions within their bicone geometry (e.g., Kakkad23), their model has a major advantage in reproducing various substructures, showing excellent agreements with the observations of both nearby and high-z AGNs (Cresci23; Marconcini23; Marconcini25a; Marconcini25b).
+Since outflows often exhibit clumpy regions within their bicone geometry (e.g., Kakkad et al., 2023), their model has a major advantage in reproducing various substructures, showing excellent agreements with the observations of both nearby and high-z AGNs (Cresci et al., 2023; Marconcini et al., 2023, 2025a, 2025b).
 Compared to the previous models, which are mainly focused on fitting 3D data cubes of individual AGNs, our model is better suited for investigating the distributions of outflow parameters for a large sample, including the effect of seeing size and the disk component.
-Note that our model can also be utilized for reproducing the observation of individual AGNs (e.g., Shin19).
+Note that our model can also be utilized for reproducing the observation of individual AGNs (e.g., Shin et al., 2019).
 Thus, our analysis with the modified model is complementary to the previous outflow models.
-5.2 Constraints on the AGN-driven outflows
-The launching velocity is the key parameter for quantifying outflow energetics (see Section 5.4), while the opening angle is related to how strongly gas outflows interact with the ISM.
+V.2 Constraints on the AGN-driven outflows
+The launching velocity is the key parameter for quantifying outflow energetics (see Section V.4), while the opening angle is related to how strongly gas outflows interact with the ISM.
 Thus, it is essential to constrain the launching velocity and opening angle of outflows for a large sample to investigate AGN feedback.
-We report launching velocities of $\lesssim$ 1500 km s-1  and typical opening angles of 30-40 (Section 4.2).
+We report launching velocities of $\lesssim$ 1500 km s-1 and typical opening angles of 30-40 (Section IV.2).
 For a consistency check, we discuss the previous estimations of these two parameters, mostly from nearby AGNs.
-For example, Crenshaw00a; Crenshaw00b reported the inner/outer opening angles as 26/40 and the maximum outflow velocity as 1300 km s-1for NGC 1068, and inner/outer opening angles as 20/36 and the maximum outflow velocity as 750 km s-1for NGC 4151 (see also Das05; Das06).
-After including dust extinction, the outer opening angles are slightly changed to 40 for NGC 1068 and 33 for NGC 4151 (Crenshaw10b).
-Their following studies also reported outflow parameters of other nearby AGNs (Crenshaw10a; Fischer10; Fischer11; Meena21; Falcone24).
-In particular, among the cases of 17 Seyfert galaxies investigated by Fischer13, the maximum velocity agrees with our range of $V_{max}\leq 1500$km s-1, except for NGC 1068.
+For example, Crenshaw and Kraemer (2000); Crenshaw et al.
+(2000) reported the inner/outer opening angles as 26/40 and the maximum outflow velocity as 1300 km s-1 for NGC 1068, and inner/outer opening angles as 20/36 and the maximum outflow velocity as 750 km s-1 for NGC 4151 (see also Das et al., 2005, 2006).
+After including dust extinction, the outer opening angles are slightly changed to 40 for NGC 1068 and 33 for NGC 4151 (Crenshaw et al., 2010b).
+Their following studies also reported outflow parameters of other nearby AGNs (Crenshaw et al., 2010a; Fischer et al., 2010, 2011; Meena et al., 2021; Falcone et al., 2024).
+In particular, among the cases of 17 Seyfert galaxies investigated by Fischer et al.
+(2013), the maximum velocity agrees with our range of $V_{max}\leq 1500$km s-1, except for NGC 1068.
 On the other hand, the outer opening angle distribution is slightly different.
@@
 The difference can be caused by a couple of effects.
-First, AGNs with higher luminosity and stronger outflows are likely to have wider opening angles (Bae16).
+First, AGNs with higher luminosity and stronger outflows are likely to have wider opening angles (Bae and Woo, 2016).
 For example, only 13% of our type 2 AGN sample have bolometric luminosity higher than $10^{44.5}$ erg/s, while more than half of their sample (9/17) are more luminous than $10^{44.5}$ erg/s.
@@
 […]
-The maximum velocity and the outer opening angle estimated by Muller-Sanchez11; Muller-Sanchez16 are also consistent with our constraints.
-On the other hand, for the best-fit parameters from Nevin18, while the maximum velocities ($V_{max}=230-720$ km s-1) also agree with our study, most of the opening angle values are much higher than the typical range of our constraint.
+The maximum velocity and the outer opening angle estimated by Müller-Sánchez et al.
+(2011, 2016) are also consistent with our constraints.
+On the other hand, for the best-fit parameters from Nevin et al.
+(2018), while the maximum velocities ($V_{max}=230-720$ km s-1) also agree with our study, most of the opening angle values are much higher than the typical range of our constraint.
 This discrepancy also appears to be due to different biconical model and/or the characteristics of their sample (i.e., AGNs with double-peaked narrow emission lines).
-Other studies that constrained the opening angles of individual galaxies (e.g., Baron18; Riffel21; Costa-Souza24) are in agreement with our constraint.
-5.3 True Outflow Size
+Other studies that constrained the opening angles of individual galaxies (e.g., Baron et al., 2018; Riffel, 2021; Costa-Souza et al., 2024) are in agreement with our constraint.
+V.3 True Outflow Size
 […]
-In Section 4.3, we verified that the outflow size measured based on velocity widths can be significantly overestimated by the seeing, especially when a true outflow size is comparable to or smaller than the seeing size.
+In Section IV.3, we verified that the outflow size measured based on velocity widths can be significantly overestimated by the seeing, especially when a true outflow size is comparable to or smaller than the seeing size.
 To properly constrain the spatial scale of outflows, it is essential to investigate the seeing-limited cases in detail.
-Previous studies reported notable spatial features in luminous type 2 AGNs, i.e., circularly extended NLR morphologies or enhanced W80 values with flat radial profiles (e.g., Liu13a; Liu13b; Harrison14).
+Previous studies reported notable spatial features in luminous type 2 AGNs, i.e., circularly extended NLR morphologies or enhanced W80 values with flat radial profiles (e.g., Liu et al., 2013a, b; Harrison et al., 2014).
 Interestingly, our model reproduces a similar flat W80 profile when the outflow is compact and the disk emission is much weaker than the outflow.
@@
 Thus, given that their samples consist of luminous type 2 quasars, implying relatively weak disk emission, the flat W80 profiles may indicate smaller outflow sizes.
-Furthermore, Villar-Martin16 found that only 2 out of 15 AGNs in their sample show clearly resolved outflows by comparing the spatial profiles of seeing with the [O III] wing emission.
+Furthermore, Villar-Martín et al.
+(2016) found that only 2 out of 15 AGNs in their sample show clearly resolved outflows by comparing the spatial profiles of seeing with the [O III] wing emission.
 Based on spectroastrometry and seeing sizes, they constrained lower and upper limits of the outflow sizes as several hundred pc and a few kpc, respectively.
@@
 […]
-In our previous IFU studies (Karouzos16a; Bae17; Kang18; Luo21; Kim23), the outflow sizes were determined based on the $\sigma_{\mathrm{[OIII]}}$-based method.
+In our previous IFU studies (Karouzos et al., 2016; Bae et al., 2017; Kang and Woo, 2018; Luo et al., 2021; Kim et al., 2023), the outflow sizes were determined based on the $\sigma_{\mathrm{[OIII]}}$-based method.
 Our main conclusion was that the outflow sizes are mostly less than several kpc scales, indicating a lack of galactic scale feedback.
@@
 For this work, we subtracted the HWHM of the seeing size from the outflow size in quadrature to correct for the seeing effect.
-However, this correction may not be sufficient as shown in Section 4.3.
+However, this correction may not be sufficient as shown in Section IV.3.
 Thus, our measurements can be conservatively considered as upper limits, particularly for higher redshift targets among our sample ($z>0.1$).
 To test the reliability of the size measurements, we apply the $\sigma_{\mathrm{[OIII]}}$-based method to various mock data.
-For this test, we set the redshift z=0.1 as a representative case and restrict $V_{max}$ from 750 to 2000 km s-1 to focus on strong outflows (see Section 4.2).
+For this test, we set the redshift z=0.1 as a representative case and restrict $V_{max}$ from 750 to 2000 km s-1 to focus on strong outflows (see Section IV.2).
 In Figure 13, we present the measured outflow size as a function of the input (true) size.
@@
 Therefore, the $\sigma_{\mathrm{[OIII]}}$-based method provides a good size estimation within 0.3 dex, or at least can be used as a reliable upper limit, depending on the relative seeing size.
-We emphasize that although the $\sigma_{\mathrm{[OIII]}}$-based method overestimates the sizes of compact outflows, the main conclusion of Kim23, i.e., lack of global feedback, remains valid.
+We emphasize that although the $\sigma_{\mathrm{[OIII]}}$-based method overestimates the sizes of compact outflows, the main conclusion of Kim et al.
+(2023), i.e., lack of global feedback, remains valid.
 Since the true size can be more compact than the measured one, the upper limit of the size still provides strong constraints on the spatial scales of AGN feedback.
@@
 Thus, if the launching velocity is fast enough, the $\sigma_{\mathrm{[OIII]}}$-based size has no significant dependence on other parameters.
-5.4 Outflow Energetics
+V.4 Outflow Energetics
 […]
-We estimate mass outflow rates ($\dot{\mathrm{M}}_{\mathrm{out}}$), momentum rates ($\dot{\mathrm{P}}_{\mathrm{out}}$), and kinetic powers ($\dot{\mathrm{E}}_{\mathrm{out}}$) of local type 2 AGNs to investigate feedback efficiency.
-These outflow rates can be determined for conical outflows as (e.g., Maiolino12; Fiore17)
-$\dot{\mathrm{M}}_{\mathrm{out}}=3\;\mathrm{M}_{\mathrm{out}}\frac{v_{\mathrm{out}}}{\mathrm{R}_{\mathrm{out}}}$ (9)
-$\dot{\mathrm{P}}_{\mathrm{out}}=\dot{\mathrm{M}}_{\mathrm{out}}v_{\mathrm{out}}$ (10)
-$\dot{\mathrm{E}}_{\mathrm{out}}=\frac{1}{2}\dot{\mathrm{M}}_{\mathrm{out}}v^{2}_{\mathrm{out}}$ (11)
-where $\mathrm{M}_{\mathrm{out}}$, $v_{\mathrm{out}}$, and $\mathrm{R}_{\mathrm{out}}$ are ionized gas mass, outflow velocity, and outflow size, respectively.
-We estimate ionized gas mass from the extinction-corrected [O III] luminosity, using Eq.
-29 in Veilleux20 and a fixed electron density of 1000 cm-3.
-In the case of outflow velocity, various definitions were adopted by previous studies, i.e., W80 (e.g., Wylezalek20; Ruschel-Dutra21), $\sqrt{v^{2}_{\mathrm{[OIII]}}+\sigma^{2}_{\mathrm{[OIII]}}}$ (e.g., Rakshit18; Kim23), or $v_{\mathrm{broad}}+2\sigma_{\mathrm{broad}}$ (e.g., Rupke13; Fiore17) to represent the flux-weighted mean outflow velocity.
-We derive the mass-weighted mean outflow velocity, using the radial profiles of outflow velocity $v(R)=V_{max}(1-R/R_{out})$ and flux $f(R)=f_{0}e^{-5R/R_{out}}$ in our model and assuming an electron density profile, $n_{e}(R)$ $\propto$ $1/R^{2}$.
-Since the mass is proportional to L/[$_{\mathrm{[OIII]}}/n_{e}$, the mass-weighted mean velocity is derived as:
-$\begin{split}v_{\mathrm{out}}&=\frac{\int^{R_{out}}_{0}v(R)\mathrm{M}_{\mathrm{out}}(R)\,R^{2}dR}{\int^{R_{out}}_{0}\mathrm{M}_{\mathrm{out}}(R)\,R^{2}dR}\\ &=\frac{\int^{R_{out}}_{0}V_{max}(1-\frac{R}{R_{out}})\times f_{0}e^{-\frac{5R}{R_{out}}}n^{-1}_{e}(R)R^{2}\,dR}{\int^{R_{out}}_{0}f_{0}e^{-\frac{5R}{R_{out}}}n^{-1}_{e}(R)R^{2}\,dR}\\ &=V_{max}\frac{\int^{1}_{0}t^{4}(1-t)e^{-5t}\,dt}{\int^{1}_{0}t^{4}e^{-5t}\,dt}\simeq 0.31\;V_{max}\end{split}$ (12)
-The mass-weighted mean outflow velocity is obtained from the maximum velocity as $v_{\mathrm{out}}=0.31V_{max}$.
-In addition, since the momentum rate and kinetic power are proportional to $v_{\mathrm{out}}^{2}$ and $v_{\mathrm{out}}^{3}$, respectively, we calculate the mass-weighted means of $v^{2}$ and $v^{3}$ as $\sim 0.14\;V^{2}_{max}$ and $\sim 0.07\;V^{3}_{max}$.
-These relations are then applied to calculate the three outflow rate parameters.
+In this section, we present estimates of the mass outflow rate ($\dot{\mathrm{M}}_{\mathrm{out}}$), momentum rate ($\dot{\mathrm{P}}_{\mathrm{out}}$), and kinetic power ($\dot{\mathrm{E}}_{\mathrm{out}}$) for local type 2 AGNs to assess the efficiency of AGN feedback.
+For conical outflows, these quantities are computed as (e.g., Maiolino et al., 2012; Fiore et al., 2017)
+$\dot{\mathrm{M}}_{\mathrm{out}}=3\,\mathrm{M}_{\mathrm{out}}\frac{v_{\mathrm{out}}}{\mathrm{R}_{\mathrm{out}}},$ (9)
+$\dot{\mathrm{P}}_{\mathrm{out}}=\dot{\mathrm{M}}_{\mathrm{out}}\,v_{\mathrm{out}},$ (10)
+$\dot{\mathrm{E}}_{\mathrm{out}}=\frac{1}{2}\,\dot{\mathrm{M}}_{\mathrm{out}}\,v_{\mathrm{out}}^{2},$ (11)
+[…]
+A variety of velocity definitions have been used in previous studies to approximate the characteristic outflow velocity, including W80 (e.g., Wylezalek et al., 2020; Ruschel-Dutra et al., 2021), $\sqrt{v_{\rm[OIII]}^{2}+\sigma_{\rm[OIII]}^{2}}$ (e.g., Rakshit and Woo, 2018; Kim et al., 2023), and $v_{\rm broad}+2\sigma_{\rm broad}$ (e.g., Rupke and Veilleux, 2013; Fiore et al., 2017).
+In our analysis, we derive a mass-weighted mean outflow velocity based on the radial profiles of the intrinsic velocity ($v(R)=V_{\rm max}(1-R/R_{\rm out})$) and flux ($f(R)=f_{0}e^{-5R/R_{\rm out}}$) in our model, assuming an electron density profile $n_{e}(R)\propto R^{-2}$.
+Since the mass is proportional to L/[$_{\mathrm{[OIII]}}/n_{e}$, the mass-weighted mean velocity becomes
+$\begin{split}v_{\mathrm{out}}&=\frac{\int_{0}^{R_{\rm out}}v(R)\,\mathrm{M}_{\mathrm{out}}(R)\,R^{2}\,dR}{\int_{0}^{R_{\rm out}}\mathrm{M}_{\mathrm{out}}(R)\,R^{2}\,dR}\\ &=\frac{\int_{0}^{R_{\rm out}}V_{\rm max}\!\left(1-\frac{R}{R_{\rm out}}\right)f_{0}e^{-5R/R_{\rm out}}n_{e}^{-1}(R)\,R^{2}\,dR}{\int_{0}^{R_{\rm out}}f_{0}e^{-5R/R_{\rm out}}n_{e}^{-1}(R)\,R^{2}\,dR}\\ &=V_{\rm max}\,\frac{\int_{0}^{1}t^{4}(1-t)e^{-5t}\,dt}{\int_{0}^{1}t^{4}e^{-5t}\,dt}\simeq 0.31\,V_{\rm max}.\end{split}$ (12)
+Thus, the mass-weighted mean outflow velocity is approximately $v_{\mathrm{out}}=0.31\,V_{\rm max}$.
+Since the momentum and kinetic power scale with $v_{\mathrm{out}}^{2}$ and $v_{\mathrm{out}}^{3}$, respectively, their mass-weighted mean values correspond to $\sim 0.14\,V_{\rm max}^{2}$ and $\sim 0.07\,V_{\rm max}^{3}$.
+These relations are used to calculate $\dot{\mathrm{M}}_{\mathrm{out}}$, $\dot{\mathrm{P}}_{\mathrm{out}}$, and $\dot{\mathrm{E}}_{\mathrm{out}}$.
 Note that the launching velocity cannot be directly measured from observational data.
-Instead, we test potential proxies of $V_{max}$ that can be calculated from the observed V[OIII] and $\sigma_{\mathrm{[OIII]}}$, using the mock SDSS data (Section 4.2).
-We find that two proxies show good agreement with the input $V_{max}$.
-One is a quadratic sum of $|\mathrm{V}_{\mathrm{[OIII]}}|$ and 3 $\sigma_{\mathrm{[OIII]}}$, which is similar to Bae16 or RodriguezZaurin13, and the other is a direct summation ($|\mathrm{V}_{\mathrm{[OIII]}}|$ + 3 $\sigma_{\mathrm{[OIII]}}$), which is similar to Rupke13 or Fiore17.
-In Figure 14, we compare the two proxies, showing that both reproduce the launching velocity with a scatter of 0.2 dex.
-Here, we use the first proxy (i.e., quadratic sum).
-Lastly, we utilize two outflow sizes as upper and lower limits.
-As discussed in Section 5.3, we first adopt our $\sigma_{\mathrm{[OIII]}}$-based sizes calculated from the kinematic outflow size–luminosity relation as upper limits and determine the lower limits of outflow rates.
-Second, we utilize the NLR sizes calculated from the NLR size–luminosity relation, which was derived from nearby AGNs (Polack24).
-We determine lower limits of the outflow sizes by dividing the NLR sizes by three, which is a typical Rout/R[OIII] ratio from Fischer18; Polack24, and calculate upper limits of the outflow rates.
-Note that these outflow sizes are typically 0.4 - 0.6 dex smaller than the $\sigma_{\mathrm{[OIII]}}$-based sizes within the dynamic range of the [O III] luminosity.
-Figure 15: Mass outflow rate (log $\dot{\mathrm{M}}_{\mathrm{out}}$, top), momentum rate (log $\dot{\mathrm{P}}_{\mathrm{out}}$, middle), and kinetic power (log $\dot{\mathrm{E}}_{\mathrm{out}}$, bottom) as a function of bolometric luminosity of the SDSS type 2 AGNs.
+Instead, we test several potential proxies for $V_{\rm max}$ that can be computed from the observed V[OIII] and $\sigma_{\rm[OIII]}$, using the mock SDSS sample (Section IV.2).
+We find that two commonly used prescriptions show good agreement with the input $V_{\rm max}$.
+The first is the quadratic sum of $|\mathrm{V}_{\rm[OIII]}|$ and $3\,\sigma_{\rm[OIII]}$, similar to that adopted by Bae and Woo (2016) and Rodríguez Zaurín et al.
+(2013).
+The second is the direct sum $|\mathrm{V}_{\rm[OIII]}|+3\,\sigma_{\rm[OIII]}$, following Rupke and Veilleux (2013) and Fiore et al.
+(2017).
+Figure 14 demonstrates that both proxies recover the launching velocity with a scatter of $\sim$0.2 dex.
+In the following analysis, we adopt the first proxy (i.e., the quadratic sum).
+Lastly, we use two different outflow size estimates to set upper and lower limits on the energetics.
+First, we adopt the $\sigma_{\rm[OIII]}$-based sizes inferred from the kinematic outflow size–luminosity relation as upper limits (see Section V.3), and compute the corresponding lower limits on the outflow rates.
+Second, we employ the NLR sizes from the NLR size–luminosity relation derived for nearby AGNs (Polack et al., 2024).
+We convert these NLR sizes to outflow sizes by dividing by three, following the typical ratio of $\mathrm{R_{out}}/\mathrm{R_{[OIII]}}$ reported in Fischer et al.
+(2018) and Polack et al.
+(2024), and then calculate the corresponding upper limits on the outflow rates.
+These NLR-based sizes are typically 0.4–0.6 dex smaller than the $\sigma_{\rm[OIII]}$-based sizes in the dynamic range of the [O III] luminosity.
+Figure 15: Mass outflow rate (log $\dot{\mathrm{M}}_{\mathrm{out}}$, top), momentum rate (log $\dot{\mathrm{P}}_{\mathrm{out}}$, middle), and kinetic power (log $\dot{\mathrm{E}}_{\mathrm{out}}$, bottom) as a function of the bolometric luminosity of the SDSS type 2 AGNs.
 The red and blue contours show the energetics based on the upper limits (red contours) and lower limits (blue contours) of the outflow sizes.
 The energetics of individual AGNs from the literature are denoted with various symbols.
-In Figure 15, we compare the upper and lower limits of the outflow rates with luminosity to investigate whether the outflows are energetic enough for efficient feedback.
-Despite the large scatter, the estimated outflow rates (contours) fall within the ranges of $\dot{\mathrm{M}}_{\mathrm{out}}\sim 1-100\;\mathrm{L}_{\mathrm{bol}}/\eta c^{2}\;(\eta=0.1)$, $\dot{\mathrm{P}}_{\mathrm{out}}\sim 0.01-1\;\mathrm{L}_{\mathrm{bol}}/c$, and $\dot{\mathrm{E}}_{\mathrm{out}}\sim 10^{-6}-10^{-3}\;\mathrm{L}_{\mathrm{bol}}$.
-These ranges are consistent with previous studies, except for some discrepancies due to different definitions of outflow parameters.
-For example, Fiore17 sample, which has kinetic powers mostly in the range of 0.01 - 1% of Lbol, has much higher $v_{\mathrm{out}}$ values since it was defined based on the kinematics of the broad component ($|v_{\mathrm{broad}}|+2\sigma_{\mathrm{broad}}$).
-In contrast, we compute $v_{\mathrm{out}}$ using the first and second moments of the total [O III] profile, leading to smaller $v_{\mathrm{out}}$ at a given luminosity.
-This difference becomes more evident in $\dot{\mathrm{E}}_{\mathrm{out}}$ since it is proportional to $v^{3}_{\mathrm{out}}$.
-Theoretical studies have suggested thresholds of kinetic power relative to bolometric luminosity (i.e., $\dot{\mathrm{E}}_{\mathrm{out}}/\mathrm{L}_{\mathrm{bol}}\sim 0.5-5\%$, e.g., Hopkins10a; Zubovas12; King15), that can suppress star formation.
-The estimated kinetic power is mostly much lower than the thresholds, implying that feedback may not be efficient.
-However, it is important to note that these outflow rates are based only on ionized outflows, which contribute a relatively low fraction of the total outflow mass (e.g., Fluetsch19).
-Nevertheless, 0.85% or 0.14% of the sample has $\dot{\mathrm{E}}_{\mathrm{out}}/\mathrm{L}_{\mathrm{bol}}$ higher than 0.5% when using outflow sizes based on Polack24 or Kim23 relations, respectively.
-These AGNs with extremely high kinetic powers also exhibit high outflow velocities, making them excellent targets for spatially resolved outflow studies in detail.
-5.5 Limitations of the current model
-Despite the modifications, there are still limitations in the model and the mock data analysis.
-First, since we adopt analytic functions for the flux and velocity distributions of outflows, our model cannot describe the detailed structures in outflows.
-For example, various studies of nearby AGNs found clumpy regions within the bicone structure (e.g., Venturi18; Mingozzi19; Kakkad23).
-Second, our model simulates no flux errors in the mock data.
-For example, the $\sigma_{\mathrm{[OIII]}}$-based method adopts the radius of the furthest [O III]-detected pixel as an outflow size when the $\sigma_{\mathrm{[OIII]}}/\sigma_{*}$ does not converge to unity.
-In this case, the S/N becomes critical for determining the outflow size.
-Also, we did not test spectroastrometry-based methods (e.g., Carniani15; Singha22), which define the outflow size as a spatial offset between the center and flux centroid of the [O III] emission.
-Reasonable errors are necessary to fit the 2D flux map when estimating the flux centroid.
-Third, we assumed that all PDFs in our MC simulations are independent since we do not have a well-described dependence among them.
-However, since some of the input parameters are likely to be correlated, the PDFs should be described as multivariate.
-Previous studies investigated the correlations between outflow parameters and bolometric luminosity (or Eddington ratio).
-The opening angle and outflow velocity are also suggested to have positive (Bae16), negative (Muller-Sanchez11), or no correlations (Fischer13).
-Therefore, MC simulations with a proper treatment of correlations between outflow parameters may better reproduce the observed gas kinematics.
-6 Summary
+[…]
+[…]
+V.5 Limitations of the current model
+Despite these improvements, our model and mock data analysis still have several limitations.
+First, because we adopt analytic functions for the flux and velocity distributions, the model cannot capture the detailed substructures of real outflows.
+High-resolution studies of nearby AGNs have revealed clumpy and irregular features within the bicone (e.g., Venturi et al., 2018; Mingozzi et al., 2019; Kakkad et al., 2023), which are not represented in our simplified prescription.
+Second, our mock data do not include flux uncertainties.
+For instance, the $\sigma_{\rm[OIII]}$-based method defines the outflow size as the radius of the outermost [O III]-detected pixel when the $\sigma_{\mathrm{[OIII]}}/\sigma_{*}$ does not converge to unity; in such cases, the S/N becomes critical in determining the size.
+We also did not test spectroastrometry-based techniques (e.g., Carniani et al., 2015; Singha et al., 2022), which estimate the outflow size from the spatial offset between the center and the flux centroid of the [O III] emission.
+These methods require realistic flux errors to properly fit the 2D flux distribution and determine the centroid.
+Third, we assume that all parameter PDFs in the MC simulations are independent, owing to the lack of well-established relations among them.
+In reality, several outflow parameters are expected to be correlated, implying that multivariate PDFs may be more appropriate.
+Previous studies have reported correlations between outflow properties and bolometric luminosity (or Eddington ratio).
+The opening angle and outflow velocity are also suggested to have positive (Bae and Woo, 2016), negative (Müller-Sánchez et al., 2011), or no correlations (Fischer et al., 2013).
+Future MC simulations incorporating such parameter correlations may better reproduce the observed gas kinematics.
+VI Summary
 In this study, we present the modified outflow model and its applications to local type 2 AGNs. Here we summarize the main results of this study.
@@
 Nevertheless, this still supports the lack of global feedback since the previously reported sizes can be considered as upper limits.
-This work has been supported by the Basic Science Research Program through the National Research Foundation of Korean Government (2021R1A2C3008486).
+This work has been supported by the Basic Science Research Program through the National Research Foundation of the Korean Government (2021R1A2C3008486).